### v1

# Random attractors on countable state spaces

Abstract
We study the synchronization behavior of discrete-time Markov chains on countable state spaces. Representing a Markov chain in terms of a random dynamical system, which describes the collective dynamics of trajectories driven by the same noise, allows for the characterization of synchronization via random attractors.
We establish the existence and uniqueness of a random attractor under mild conditions and show that forward and pullback attraction are equivalent in our setting. Additionally, we provide necessary and sufficient conditions for reaching the random attractor, or synchronization respectively, in a time of finite mean.
By introducing insulated and synchronizing sets, we structure the state space with respect to the synchronization behavior and characterize the shape of the random attractor.
1 Introduction
In this work, we consider discrete-time Markov chains on a countable state space which are irreducible and positive recurrent, and, hence, admit a unique stationary distribution. The law of a Markov chain $(X_{n})_{n\in\mathds{N}}$ is governed by the transition probabilities $\mathds{P}(X_{n+1}=y\mid X_{n}=x)$. These probabilities describe the random motion of a single particle, but contain no information about how multiple particles move collectively, i.e. how the individual trajectories are coupled. We consider a specific type of coupling, which arises from representing a Markov chain by a random dynamical system, short RDS, [Arn98].
Random dynamical systems model in a pathwise manner the evolution of stochastic processes for fixed noise realizations, allowing the simultaneous tracking of multiple trajectories under the same randomness. RDS representations of Markov chains on discrete state spaces have previously been studied in [YWQ16, GKM21, EOMUW24]. Informally, an RDS representation of a Markov chain randomly picks a map $f_{n}:\mathds{X}\to\mathds{X}$ in each time step $n\in\mathds{N}$ such that the process $X_{n+1}=f_{n}(X_{n})$ admits the law of the given Markov chain. For a formal definition of RDS representations of a Markov chain, we refer to Section 2. Note that such a representation naturally describes the collective behavior of any number of particles, i.e. $Y_{n+1}=f_{n}(Y_{n})$ for another starting condition $Y_{0}$. Note that if $X_{n}=Y_{n}$ for some $n\in\mathds{N}$, then $X_{m}=Y_{m}$ for all $m\geq n$. This is what we call synchronization. If any two initial conditions $X_{0}$, $Y_{0}$ eventually synchronize, we call the RDS synchronizing. We formalize the synchronization behavior of trajectories using the notion of a random attractor.
Related work
For RDS on general state spaces, the theory of attractors has been thoroughly developed over the past decades, see for instance [CF94, CDF97, Arn98, CK15, Cra16]. Generally, different notions of attraction coexist without redundancy, where examples have been typically studied in the context of stochastic differential equations [Sch02, CS04, CDS09, DS11]. In more detail: as opposed to the autonomous deterministic setting, random dynamical systems (being non-autonomous) may exhibit different notions of attraction, e.g. pullback and forward attractors [KR11]. Additionally, random attractors are random sets whose notion of attraction is further classified by the type of convergence, i.e. in the strong or the weak sense [CK15]. One can also distinguish random attractors in terms of the family of subsets which are attracted towards them. We show that in our setting, which considers discrete states space, almost all of the commonly used notions of attraction are equivalent, cf. Theorem 3.1 and Remark 3.2 thereafter. This is a direct generalization of results formulated in [EOMUW24].
The representation of a Markov chain by an RDS is very similar to the notion of a stochastic recursive sequence [BF92]. More recently, the representation of Markov chains by stochastic recursive sequences has been studied using the framework of unimodular networks [BHMK18, BHMM19]. One of the key motivations for these works is the coupling from the past algorithm, which was introduced in [PW96] and provides a method to draw perfect samples from the stationary distribution of a Markov chain. The objects we consider in this work are similar to those studied in the references mentioned above, despite being formulated in a different mathematical language. For example, a random attractor corresponds to a finite set of bi-infinite paths in [BHMM19]. Using the framework of random dynamical systems, which enables the use of powerful tools like the correspondence theorem, cf. Theorem 2.7, we obtain the existence of a unique random attractor, characterize its size, and derive a detailed structuring of the state space into insulated and synchronizing sets, cf. Section 4. Furthermore, Section 5 concerns the expected time until attraction, or synchronization respectively. This quantity has previously been considered in [BHMM19]. We provide necessary and sufficient conditions for the expected hitting times to be finite.
Main results and structure of the paper
The main results of our work are the following.
  • Theorem 3.1. There exists a unique random attractor that is both a pullback attractor for finite deterministic sets and a forward attractor for compact random sets. This shows that for finite deterministic sets forward and pullback attraction are equivalent.
  • Theorem 5.2. The expected time to hit the attractor is finite for every initial condition if and only if the associated Markov chain is ergodic of degree 2.
  • Theorem 5.4. For synchronizing RDS, the expected time until two given initial conditions synchronize is always finite.
In Section 2, we recall some theory of Markov chains on countable state spaces and formally introduce random dynamical systems and different notions of random attractors. Section 3 is dedicated to the proof of Theorem 3.1 and establishes the existence of a unique random attractor $A$ that has a fixed size $\kappa$. In Section 4 we structure the state space with respect to the synchronizing behavior and introduce insulated and synchronizing sets. In particular, we show that the random attractor $A$ is a maximum insulated set (Theorem 4.5), and derive restrictions to the possible values of $\kappa$ (see also Remark 4.8). In Section 5, we study how quickly initial conditions hit the attractor, as well as how quickly synchronization occurs in the case of synchronizing RDS. In particular, we prove Theorem 5.2 and Theorem 5.4, which provide finite expected times until attraction, or synchronization respectively. Finally, we discuss our findings in Section 6 where we also address some open questions.
2 Preliminaries
2.1 Markov chains
We consider time-homogeneous discrete-time Markov chains $(X_{n})_{n\in\mathds{N}}$ on a countable state space $\mathds{X}$. A Markov chain is defined by its transition probabilities for each pair $x,y\in\mathds{X}$
$P(x,y):=\mathds{P}(X_{1}=y\>|\>X_{0}=x),\quad P^{n}(x,y):=\mathds{P}(X_{n}=y\>|\>X_{0}=x).$ (1)
A stationary distribution of $(X_{n})_{n\in\mathds{N}}$ is a probability vector $\pi\in\mathds{R}^{\mathds{X}}_{\geq 0}$ that satisfies
$(\pi P)(y):=\sum_{x\in\mathds{X}}\pi(x)P(x,y)=\pi(y),\quad\forall y\in\mathds{X},\quad\text{and}\quad\sum_{x\in\mathds{X}}\pi(x)=1.$
We call a Markov chain irreducible if for each $x,y\in\mathds{X}$ there is an $n\in\mathds{N}$ such that $P^{n}(x,y)>0$. An irreducibility class $\mathcal{I}\subset\mathds{X}$ is a set such that the Markov chain restricted to $\mathcal{I}$ is irreducible, i.e. for all $x,y\in\mathcal{I}$ there is an $n\in\mathds{N}$ such that $P^{n}(x,y)$, and for all $x\in\mathcal{I}$ and $y\in\mathds{X}\setminus\mathcal{I}$ we find $P^{n}(x,y)=0$ for all $n\in\mathds{N}$.
Denote the first hitting time of a state $x\in\mathds{X}$ by
$\tau_{x}:=\min\{n\geq 1\>|\>X_{n}=x\}.$ (2)
We denote the probability and the expectation, conditioned on $X_{0}=x$, by $\mathds{P}_{x}$ and $\mathds{E}_{x}$ respectively. Hence, the expected return time of a state $x\in\mathds{X}$ is written as $\mathds{E}_{x}[\tau_{x}]$. We call the Markov chain recurrent if
$\mathds{P}_{x}(\tau_{x}<\infty)=1,\quad\forall x\in\mathds{X}.$
We call the Markov chain positive recurrent if the expected return times are of finite mean for all states, i.e.
$\mathds{E}_{x}[\tau_{x}]<\infty,\quad\forall x\in\mathds{X}.$
The following equivalence is well-known, see e.g. [BH22, Theorem 2.6].
Theorem 2.1.
For an irreducible Markov chain $(X_{n})_{n\in\mathds{N}}$ on a countable state space, the following notions are equivalent:
  1. (i)
    There is an $x\in\mathds{X}$ such that $\mathds{E}_{x}[\tau_{x}]<\infty$;
  2. (ii)
    $(X_{n})_{n\in\mathds{N}}$ is positive recurrent;
  3. (iii)
    $(X_{n})_{n\in\mathds{N}}$ admits a unique stationary distribution $\pi$ that satisfies $\pi(x)=\mathds{E}_{x}[\tau_{x}]^{-1}>0$, for any $x\in\mathds{X}$.
Under these equivalent conditions:
  • For any $x,y\in\mathds{X}$ we find the $\mathds{P}_{x}$-a.s. limit
    $\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}\mathds{1}_{y}(X_{n})=\pi(y);$ (3)
  • $\mathds{E}_{x}[\tau_{y}]<\infty$, for all $x,y\in\mathds{X}$.
In the following, we always work with irreducible and positive recurrent Markov chains whose unique stationary distribution is denoted by $\pi$. Note that we do not assume aperiodicity.
2.2 Random dynamical systems
We briefly introduce the general notion of a random dynamical system (RDS), before focusing on the representation of a Markov chain by an RDS. Random dynamical systems model the evolution of stochastic processes pathwise for fixed noise realizations. This allows for simultaneously tracking multiple trajectories under the same randomness, giving rise to interesting phenomena like random attractors. Since we only consider countable state spaces, we directly provide all definitions in the setting of a countable space $\mathds{X}$ which we equip with the powerset $\sigma$-algebra, denoted by $\mathcal{P}(\mathds{X})$. For a detailed introduction to random dynamical systems on general state spaces in discrete and continuous time, we refer to [Arn98].
An RDS $(\theta,\varphi)$ on $\mathds{X}$ consists of a probability space $(\Omega,\mathcal{F},\mathds{P})$, a group of bimeasurable measure-preserving transformations $(\theta_{n}:\Omega\to\Omega)_{n\in\mathds{Z}}$ such that for all $\omega\in\Omega$
$\begin{split}\theta_{0}\omega&=\omega,\\ \theta_{n+m}\omega&=(\theta_{n}\circ\theta_{m})\omega,\quad\forall m,n\in\mathds{Z},\end{split}$ (4)
and a $(\mathcal{F}\otimes\mathcal{P}(\mathbb{N})\otimes\mathcal{P}(\mathds{X}))$-measurable map $\varphi:\Omega\times\mathds{N}\times\mathds{X}\to\mathds{X}$ that has the cocycle property
$\begin{split}\varphi_{\omega}^{0}(x)&=x,\\ \varphi^{n+m}_{\omega}(x)&=\left(\varphi^{n}_{\theta_{m}\omega}\circ\varphi^{m}_{\omega}\right)(x),\end{split}$ (5)
for all $\omega\in\Omega$, $m,n\in\mathds{N}$ and $x\in\mathds{X}$. Due to the cocycle property, an RDS is uniquely defined by the time-one maps $\varphi_{\omega}^{1}:\mathds{X}\to\mathds{X}$ and $\theta_{1}:\Omega\to\Omega$. We usually write $\varphi_{\omega}$ and $\theta$ instead of $\varphi_{\omega}^{1}$ and $\theta_{1}$. We additionally assume that $\theta$ is ergodic with respect to $\mathds{P}$, which implies that $\theta_{-1}$ is ergodic with respect to $\mathds{P}$ as well. From the cocycle property it follows that if two initial states $x,y\in\mathds{X}$ coalesce, i.e. $\varphi_{\omega}^{n}(x)=\varphi_{\omega}^{n}(y)$ for some $n\in\mathds{N}$, then $\varphi_{\omega}^{m}(x)=\varphi_{\omega}^{m}(y)$ for all $m>n$. We say that $x$ and $y$ $\omega$-synchronize [HQW+20].
The maps $\varphi_{\omega}^{n}$ can also be applied to sets and tuples. For $K$ being a set or tuple of elements in $\mathds{X}$, we write $\varphi_{\omega}^{n}(K)$ for the set or tuple consisting of $\varphi_{\omega}^{n}(x)$, $x\in K$. Whether $\varphi^{n}_{\omega}$ is applied to an element, a set or a tuple should be apparent from the context.
Next, we define invariant measures of an RDS. The RDS $(\theta,\varphi)$ induces the so-called skew-product map on $\Omega\times\mathds{X}$ given by
$\Phi:\Omega\times\mathds{X}$ $\to\Omega\times\mathds{X},$
$(\omega,x)$ $\mapsto(\theta\omega,\varphi_{\omega}(x)).$
An invariant measure of the RDS $(\theta,\varphi)$ will be defined to be invariant with respect to this skew-product map. For any measure $\mu$ on a measurable space $M$ and a measurable map $f:M\to M$, we define the push-forward of $\mu$ by $f$,
$f_{*}\mu(\cdot)=\mu(f^{-1}(\cdot)).$
Hence, the push-forward operator $f_{*}$ maps measures on $M$ to measures of $M$. In particular, $f_{*}$ maps probability measures to probability measures.
Definition 2.2.
A probability measure $\mu$ on $(\Omega\times\mathds{X},\mathcal{F}\otimes\mathcal{P}(\mathds{X}))$ is called invariant if
  1. (i)
    $\Phi_{*}\mu=\mu$;
  2. (ii)
    The marginal of $\mu$ on $\Omega$ is $\mathds{P}$ such that $\mu$ can be factorized uniquely into
    $\mu(d\omega,x)=\mu_{\omega}(x)\mathds{P}(d\omega),$
    where the $\mu_{\omega}$ are the sample measures on $\mathds{X}$, i.e.
    • $\mu_{\omega}$ is a probability measure on $\mathds{X}$ for $\mathds{P}$-a.e. $\omega\in\Omega$;
    • $\omega\mapsto\mu_{\omega}(x)$ is measurable for any $x\in\mathds{X}$.
Invariance of a measure $\mu$ on $\Omega\times\mathds{X}$ with marginals $\mathds{P}$ can be checked directly via the invariance of the sample measures. This alternative characterization tends to be much easier to work with.
Proposition 2.3.
A probability measure $\mu$ on $(\Omega\times\mathds{X},\mathcal{F}\otimes\mathcal{P}(\mathds{X}))$ with marginal $\mathds{P}$ on $\Omega$ is invariant with respect to the RDS $(\theta,\varphi)$, if and only if the samples measures satisfy
$(\varphi_{\omega})_{*}\mu_{\omega}=\mu_{\theta\omega},$ (6)
for $\mathds{P}$-a.e. $\omega\in\Omega$.
A proof of this result can be found in [KS12, Proposition 1.3.27]. Note that due to the cocycle property, equation (6) implies that $(\varphi^{n}_{\omega})_{*}\mu_{\omega}=\mu_{\theta_{n}\omega}$ for for all $n\in\mathds{N}$ and $\mathds{P}$-a.e. $\omega\in\Omega$.
2.3 Markov RDS
Let $(\theta,\varphi)$ be an RDS on a countable state space $\mathds{X}$. For any $s,t\in\mathds{Z}$ with $s<t$, we denote by $\mathcal{F}_{s}^{t}\subset\mathcal{F}$ the $\sigma$-algebra generated by (the subsets of $\mathcal{F}$ of zero measure and) the random variables $\omega\mapsto\varphi^{n}_{\theta_{m}\omega}(x)$ for $x\in\mathds{X}$ and $m,n\in\mathds{Z}$ with $s\leq m\leq m+n\leq t$. We define the $\sigma$-algebras
$\mathcal{F}_{-\infty}^{t}$ $:=\sigma(\mathcal{F}_{s}^{t}\>|\>s\in\mathds{Z},s<t),$ (7)
$\mathcal{F}_{s}^{\infty}$ $:=\sigma(\mathcal{F}_{s}^{t}\>|\>t\in\mathds{Z},s<t),$
$\mathcal{F}_{-\infty}^{\infty}$ $:=\sigma(\mathcal{F}_{-t}^{t}\>|\ \ t\in\mathds{Z}).$
Definition 2.4.
An RDS $(\theta,\varphi)$ is called Markov if $\mathcal{F}_{-\infty}^{0}$ is independent from $\mathcal{F}_{0}^{\infty}$.
The $\sigma$-algebra $\mathcal{F}_{-\infty}^{\infty}$ contains all the dynamical information of the RDS. Clearly, $\mathcal{F}_{-\infty}^{\infty}\subset\mathcal{F}$, but in general $\mathcal{F}$ might contain strictly more sets than $\mathcal{F}_{-\infty}^{\infty}$. Note that $(\theta,\varphi)$ is a well-defined RDS over the probability space $(\Omega,\mathcal{F}_{-\infty}^{\infty},\mathds{P})$. In the Appendix A, we show that the space $(\Omega,\mathcal{F}_{-\infty}^{\infty},\mathds{P})$ is effectively a shift space and state some additional results in this setting like the indecomposability of attractors and invariant Markov measure. Since our main results do not require this assumption, we work with general $\mathcal{F}$.
Note that for a Markov RDS, $\mathcal{F}_{-\infty}^{t}$ is independent from $\mathcal{F}_{t}^{\infty}$ for any $t\in\mathds{Z}$. In particular, the $\sigma$-algebras $\mathcal{F}_{t}^{t+1}$ are mutually independent, and the maps $\varphi_{\omega}$ and $\varphi_{\theta_{n}\omega}$ are independent for all $n\in\mathds{Z}\setminus\{0\}$. These observations imply that, given a Markov RDS $(\theta,\varphi)$, the random process defined by
$X_{0}(\omega)=x,\quad X_{n}(\omega)=\varphi^{n}_{\omega}(x),$ (8)
is a time-homogeneous Markov chain [Arn98, Theorem 2.1.4]. The transition probabilities $P$ of $X$ are given by
$P(x,y)=\mathds{P}(\varphi_{\omega}(x)=y).$
We say that the RDS represents the Markov chain $(X_{n})_{n\in\mathds{N}}$. Conversely, we call $(X_{n})_{n\in\mathds{N}}$ the associated Markov chain of the RDS. The key quality of an RDS representation is that it models the simultaneous motion of arbitrarily many points in $\mathds{X}$ under the same noise realisation $\omega\in\Omega$, while the Markov chain only describes the random motion of a single point in $\mathds{X}$.
Remark 2.5.
While each Markov RDS represents a unique Markov chain, a given Markov chain with transition probabilities $P$ has, in general, many possible RDS representations, see e.g. [YWQ16]. In this work, we derive statements about random dynamical systems representing given Markov chains, independently from which representation is chosen.
For Markov RDS, we strengthen the notion of an invariant measure, cf. Definition 2.2, and introduce the notion of an invariant Markov measure.
Definition 2.6.
Let $(\theta,\varphi)$ be a Markov RDS. An invariant measure $\mu$ is called an invariant Markov measure if the samples measures $\mu_{\omega}$ are $\mathcal{F}_{-\infty}^{0}$-measurable, i.e. for each $x\in\mathds{X}$, the map $\omega\mapsto\mu_{\omega}(x)$ is $\mathcal{F}_{-\infty}^{0}$-measurable.
A central tool for our work is the correspondence theorem, which links invariant Markov measures of a Markov RDS to the stationary distributions of the associated Markov chain in a one-to-one correspondence. In particular, if the associated Markov chain has a unique stationary distribution, the representing RDS has a unique invariant Markov measure. For a proof of the following theorem, see [KS12, Theorem 4.2.9].
Theorem 2.7 (Correspondence Theorem).
Let $(\theta,\varphi)$ be a Markov RDS. The following assertions hold:
  1. (i)
    Let $\mu$ be an invariant Markov measure on $\Omega\times\mathds{X}$ with sample measures $\mu_{\omega}$. Then
    $\pi:=\mathds{E}[\mu_{\omega}]$
    is a stationary distribution of the associated Markov chain.
    Additionally, if $\mu^{\prime}$ is another invariant Markov measure with sample measures $\mu^{\prime}_{\omega}$ such that $\mathds{E}[\mu_{\omega}]=\mathds{E}[\mu^{\prime}_{\omega}]$, then $\mu=\mu^{\prime}$.
  2. (ii)
    Let $\pi$ be a stationary distribution of the associated Markov chain. Then the the weak limit
    $\mu_{\omega}:=\lim_{n\to\infty}(\varphi^{n}_{\theta_{-n}\omega})_{*}\pi,$ (9)
    exists for $\mathds{P}$-a.e. $\omega\in\Omega$ and the measure $\mu$ defined by the sample measures $\mu_{\omega}$ is an invariant Markov measure such that $\mathds{E}[\mu_{\omega}]=\pi$.
Note that the correspondence theorem only establishes a one-to-one correspondence between stationary distributions of the associated Markov chain and invariant Markov measures on $\Omega\times\mathds{X}$. In general, there can be other, non-Markovian invariant measures on $\Omega\times\mathds{X}$. Under the assumption $\mathcal{F}=\mathcal{F}_{-\infty}^{\infty}$, [Arn98, Theorem 1.7.2] shows that the invariant Markov measure given by $\mu_{\omega}=\lim_{n\to\infty}(\varphi^{n}_{\theta_{-n}\omega})_{*}\pi$ is ergodic with respect to the skew-product map $\Phi$ if and only if $\pi$ is an ergodic stationary distribution of the associated Markov chain. In particular, an RDS representing an irreducible and positive recurrent Markov chain has a single invariant measure. This measure is Markovian and ergodic with respeect to $\Phi$. We discuss this result and its implications in Appendix A and construct an example of an RDS that has more than one invariant measure.
2.4 Prototypical constructions
We introduce well-known constructions for Markov RDS representing a given Markov chain. We will use these constructions in most of the examples in the following sections.
Firstly, we introduce the shift space. Given a Markov chain on a countable state space $\mathds{X}$ with transition probabilities $P$, consider the set $\Sigma$ of all maps $f:\mathds{X}\to\mathds{X}$. We equip $\Sigma$ with the $\sigma$-algebra $\mathcal{G}$ generated by the maps $f\mapsto f(x)$, for $x\in\mathds{X}$. Let $\mathds{P}_{0}$ be a probability measure on $(\Sigma,\mathcal{G})$ satisfying $\mathds{P}_{0}(f(x)=y)=P(x,y)$ (see, for instance, equation (10) below). Define the probability space $(\Omega,\mathcal{F},\mathds{P}):=(\Sigma^{\mathds{Z}},\mathcal{G}^{\otimes\mathds{Z}},\mathds{P}_{0}^{\otimes\mathds{Z}})$. We write elements of $\Omega$ as $\omega=(\ldots,\omega_{-1},\omega_{0},\omega_{1},\ldots)$, where $\omega_{n}\in\Sigma$. It is easy to observe that the left-shift $\theta:\Omega\to\Omega$, i.e. $(\theta\omega)_{n}=\omega_{n+1}$, preserves the measure $\mathbb{P}$. Over this shift space $(\Omega,\mathcal{F},\mathds{P},\theta)$, we define a cocycle by $\varphi_{\omega}(x)=\omega_{0}(x)$ and observe that the resulting RDS $(\theta,\varphi)$ is a Markov RDS representation of the Markov chain on $\mathds{X}$ with transition probabilities $P$. The measure $\mathds{P}_{0}$ on $\Sigma$ uniquely determines the behavior of the RDS representation. In Appendix A, we show that $(\theta,\varphi)$ satisfies $\mathcal{F}=\mathcal{F}_{-\infty}^{\infty}$ and state some additional results for such RDS.
A special case of the construction above is the independent RDS. In the independent RDS points move independently from one another until they coalesce. Using Carathéodory’s extension theorem [Fol99, Theorem 1.11], define a measure $\mathds{P}_{0}$ on $(\Sigma,\mathcal{G})$ by
$\mathds{P}_{0}(f(K)=L)=\prod_{i=1}^{n}P(x_{i},y_{i}),$ (10)
where $K=(x_{1},\ldots,x_{n})$ and $L=(y_{1},\ldots,y_{n})$ are tuples of distinct elements of $\mathds{X}$. We call the resulting RDS over the shift-space $(\Sigma^{\mathds{Z}},\mathcal{G}^{\otimes\mathds{Z}},\mathds{P}_{0}^{\otimes\mathds{Z}})$ the independent RDS. We will present and analyze the independent RDS of a given Markov chain in Example 1.
2.5 Random attractors
Let $(\theta,\varphi)$ be a Markov RDS on a countable state space $\mathds{X}$. A central object of study in the theory of random dynamical systems is the notion of a random attractor. In the literature, there are multiple notions of attractors that are in general neither equivalent nor strictly stronger that one another, cf. [Sch02]. All of these definitions introduce attractors as invariant compact random sets. For a detailed discussion on the matter, we refer to [CK15]. Focusing on discrete time and a countable state space, we give the definitions of such sets accordingly, using compact and finite interchangeably.
Definition 2.8.
We call $A:\Omega\to\mathcal{P}(\mathds{X})$ a compact random set if
  1. (i)
    $\omega\mapsto\mathds{1}_{A(\omega)}(x)$ is measurable for each $x\in\mathds{X}$;
  2. (ii)
    $A(\omega)$ is compact, i.e. finite, for $\mathds{P}$-a.e. $\omega\in\Omega$.
We say that $A$ is an invariant compact random set if additionally
  1. (iii)
    $A$ is $\varphi$-invariant, i.e.
[…]
Definition 2.9.
An invariant compact random set $A$ is called a
  1. (i)
    pullback attractor for compact random sets, if for each compact random set $B$
    $\mathds{P}(\exists N\in\mathds{N}\,,\forall n\geq N:\varphi^{n}_{\theta_{-n}\omega}(B(\theta_{-n}\omega))\subset A(\omega))=1.$
  2. (ii)
    forward attractor for compact random sets, if for each compact random set $B$
    $\mathds{P}(\exists N\in\mathds{N}\,,\forall n\geq N:\varphi^{n}_{\omega}(B(\omega))\subset A(\theta_{n}\omega))=1.$
  3. (iii)
    weak attractor for compact random sets, if for each compact random set $B$
    $\lim_{n\to\infty}\mathds{P}(\varphi^{n}_{\theta_{-n}\omega}(B(\theta_{-n}\omega))\subset A(\omega))$ $=1,$ (11)
    $\lim_{n\to\infty}\mathds{P}(\varphi^{n}_{\omega}(B(\omega))\subset A(\theta_{n}\omega))$ $=1.$
  4. (iv)
    pullback/forward/weak attractor for finite deterministic sets, if the respective definition holds for finite deterministic sets $B\subset\mathds{X}$;
  5. (v)
    pullback/forward/weak point attractor, if the respective definition holds for all singleton sets $B=\{x\}$, $x\in\mathds{X}$.
Remark 2.10.
The two criteria in the definition of weak attraction are equivalent by the measure-preserving property of $\theta$. Both of these criteria are listed to emphasize that a weak attractor is both a weak forward and a weak pullback attractor.
For general RDS, pullback attraction and forward attraction are neither weaker nor stronger than the other, cf. [Sch02]. On the other hand, it follows directly from the definitions that both pullback and forward attraction are stronger than the respective weak attraction. Our results of the next section, cf. Theorem 3.1, show that in our discrete setting pullback attraction is stronger than forward attraction.
In many settings, attraction of compact random sets is too strong of a notion. Hence, it is common to restrict attraction of compact random sets to a class of ’nice’ sets. In [KR11, Definition 14.3], a pullback attractor is only required to attract so-called tempered random sets. In the next section, we show that any Markov RDS representing an irreducible and positive recurrent Markov chain admits a forward attractor for general compact random sets. Therefore, we keep the definitions at this high level of generality.
Closely related to the concept of attraction, we introduce the notion of synchronization.
Definition 2.11.
We say that $x,y\in\mathds{X}$ $\omega$-synchronize if there is an $n\in\mathds{N}$ such that $\varphi^{n}_{\omega}(x)=\varphi^{n}_{\omega}(y)$. An RDS $(\theta,\varphi)$ is called synchronizing, if any $x,y\in\mathds{X}$ $\omega$-synchronize for $\mathds{P}$-a.e. $\omega\in\Omega$.
For a synchronizing RDS, we define the synchronization time of two points $x,y\in\mathds{X}$ as
$T_{\Delta}(\omega,x,y):=\min\{n\in\mathds{N}\>|\>\varphi_{\omega}^{n}(x)=\varphi_{\omega}^{n}(y)\}.$ (12)
This time is $\mathds{P}$-a.s. finite by definition. In section 5.2, we will show the also the expected synchronization time is always finite.
If an RDS has a forward point-attractor $A$ that $\mathds{P}$-a.s. consists of a single point, i.e. $\left|A(\omega)\right|=1$ for $\mathds{P}$-a.e. $\omega\in\Omega$, then the RDS is synchronizing. For Markov RDS, representing an irreducible and positivly recurrent Markov chain, the converse holds true as well, cf. Remark 4.6. In general, the converse is not true since there are RDS that are synchronizing but do not have an attractor.
3 Existence of an attractor
The main result of this section states that an RDS representing an irreducible and positive recurrent Markov chain possesses a unique random attractor that satisfies all the notions of attraction in Definition 2.9 except pullback attraction for compact random sets. In particular, this shows that, apart from pullback attraction for compact random sets, all notions of attraction are equivalent in this setting. Therefore, in this setting, pullback attraction is a strictly stronger notion of attraction than forward attraction. We emphasize that the results of this section hold for any RDS representation of a given Markov chain.
Theorem 3.1.
Let $(\theta,\varphi)$ be a Markov RDS on a countable state space $\mathds{X}$ representing a discrete-time Markov chain which is irreducible and positive recurrent.
  1. (a)
    There is an $\mathcal{F}_{-\infty}^{0}$-measurable invariant compact random set $A$ which is a pullback attractor for finite deterministic sets and a forward attractor for compact random sets.
  2. (b)
    There is a constant $\kappa\in\mathds{N}$ such that $\left|A(\omega)\right|=\kappa$ for $\mathds{P}$-a.e. $\omega\in\Omega$.
  3. (c)
    The probability that a state $x\in\mathds{X}$ lies in the attractor is given by
    $\mathds{P}(x\in A(\omega))=\kappa\pi(x).$ (13)
The rest of this subsection is dedicated to proving this theorem. Instead of writing out one unified proof of $(a)$, we formulate multiple smaller propositions that together provide a proof. The structure is as follows:
  1. (I)
    There is a weak point attractor $A$ which is $\mathcal{F}^{0}_{-\infty}$-measurable;
  2. (II)
    $A$ is a forward point attractor;
  3. (III)
    $A$ is a pullback point attractor;
  4. (IV)
    In countable state spaces forward point attraction is equivalent to forward attraction of compact random sets and pullback point attraction is equivalent to pullback attraction for finite deterministic sets.
  5. (V)
    The statements $(b)$ and $(c)$ hold.
Remark 3.2.
Since forward attraction is strictly stronger than weak attraction, the attractor $A$ from Theorem 3.1 satisfies all notions of attraction of Definition 2.9, except pullback attraction for compact random sets. This shows that all notions, except pullback attraction for compact random sets, are equivalent. Given an invariant compact random set $A^{\prime}$ that satisfies any notion of attraction, it is at least a weak point attractor, which is the weakest type of attractor. By (IV), $A^{\prime}$ is a weak attractor for compact random sets as well. Such attractors are $\mathds{P}$-a.s. unique. Hence, $A^{\prime}$ coincides $\mathds{P}$-a.s. with $A$ from Theorem 3.1 and, hence, satisfies all notions of attraction, except pullback attraction for compact random sets.
Throughout this section, we consider a fixed RDS $(\theta,\varphi)$ representing a discrete-time Markov chain $(X_{n})_{n\in\mathds{N}}$ with transition probabilities $P$, that is irreducible and positive recurrent with stationary distribution $\pi$. We complete step (I) using results from [FGS17]. The existence of a weak attractor under strictly stronger assumptions has been shown in [EOMUW24]. The following proposition is a direct extension of this result.
Proposition 3.3 (I).
There is a weak point attractor $A$ which is $\mathcal{F}^{0}_{-\infty}$-measurable.
Proof.
This theorem uses results from [FGS17]. In their work, $\mathcal{F}=\mathcal{F}_{-\infty}^{\infty}$ is assumed. Note that $(\theta,\varphi)$ is a well-defined RDS over the probability space $(\Omega,\mathcal{F}_{-\infty}^{\infty},\mathds{P})$. Since $\mathcal{F}_{-\infty}^{\infty}\subset\mathcal{F}$, any point attractor of the RDS with $\sigma$-algebra $\mathcal{F}_{-\infty}^{\infty}$ is also a point attractor of the RDS with $\sigma$-algebra $\mathcal{F}$. Hence, w.l.o.g. we assume $\mathcal{F}=\mathcal{F}_{-\infty}^{\infty}$ whenever we use results from [FGS17] to show the existence of a point attractor.
The correspondence theorem, in particular (9), provides a unique invariant Markov measure $\mu$ for $(\theta,\varphi)$ on $\Omega\times\mathds{X}$. By [FGS17, Lemma 2.19], we know that the sample measures $\mu_{\omega}$ either $\mathds{P}$-a.s. have no point mass or $\mathds{P}$-a.s. consist of finitely many atoms of the same mass. Since we work on a countable state space that does not support probability measures without point mass, we can find a random set $A$ and $\kappa\in\mathds{N}$ such that for $\mathds{P}$-a.e. $\omega\in\Omega$
$\mu_{\omega}=\frac{1}{\kappa}\sum_{a\in A(\omega)}\delta_{a},$ (14)
where $|A(\omega)|=\kappa$. In particular, we have for such $\omega$ that
$A(\omega)=\text{supp}(\mu_{\omega}).$ (15)
Hence, $A$ is a random set that is $\mathds{P}$-a.s. compact. Note that there is a $\theta$-invariant set $\hat{\Omega}\subset\Omega$ of full $\mathds{P}$-measure such that these assertions hold for all $\omega\in\hat{\Omega}$.
To see that $A$ is $\mathds{P}$-a.s. invariant, we use the invariance of $\mu$ to find the identity
$A(\theta\omega)=\text{supp}(\mu_{\theta\omega})=\varphi_{\omega}(\text{supp}(\mu_{\omega}))=\varphi_{\omega}(A(\omega)),$ (16)
for all $\omega\in\hat{\Omega}$. This shows that $A$ is $\mathds{P}$-a.s. invariant and, in summary, an invariant compact random set. The fact that $A(\omega)$ is $\mathcal{F}_{-\infty}^{0}$-measurable follows directly from $\mu$ being a Markov measure.
The main tool to prove that $A$ is a weak point attractor is [FGS17, Proposition 2.20], which implies that $A$ is a weak point attractor for points of the set
$E_{0}:=\{x:\lim_{n\to\infty}P^{n}(x,\cdot)\rightarrow\pi\}.$ (17)
[…]
On each set $W_{i}$, we consider the time-$p$ RDS over the probability space $(\overline{\Omega},\mathcal{F},\mathds{P})$, defined via the time-$p$ map $\varphi^{p}_{\omega}:W_{i}\rightarrow W_{i}$ and the time-$p$ shift $\theta_{p}$. Define the rescaled measures of $\mu$ restricted to $W_{i}$ by
$\mu_{\omega}^{i}:=\frac{1}{\mu_{\omega}(W_{i})}\mu_{\omega}|_{W_{i}}=p\>\mu_{\omega}|_{W_{i}}.$
This measure defines an invariant Markov measure of the time-$p$ RDS on $W_{i}$ since
$(\varphi^{p}_{\omega})_{*}(\mu_{\omega}|_{W_{i}})=\mu_{\theta_{p}\omega}|_{W_{i}},$
which follows immediately from the invariance of $\mu$. Therefore, the sets defined by
$A^{i}(\omega):=\text{supp}(\mu^{i}_{\omega})=\text{supp}(\mu_{\omega})\cap W_{i},$
are invariant compact random sets that satisfy
$A^{i}(\omega)=A(\omega)\cap W_{i}.$
The associated time-$p$ Markov chain on $W_{i}$ is aperiodic, irreducible and positive recurrent. Hence, the set $E_{0}^{i}$, defined analogously to (17), equals $W_{i}$, and $A^{i}(\omega)$ is a weak point attractor of the time-$p$ RDS on $W_{i}$. For any point $x\in W_{i}$, we find
$\lim_{n\to\infty}\mathds{P}(\varphi^{pn}_{\omega}(x)\in A^{i}(\omega))=1.$
Since $A^{i}(\omega)\subset A(\omega)$ and $\varphi^{N}_{\omega}(x)\in A(\omega)$ implies $\varphi^{n}_{\omega}(x)\in A(\omega)$ for any $n\geq N$, we find
$\lim_{n\to\infty}\mathds{P}(\varphi^{n}_{\omega}(x)\in A(\omega))=1.$
This holds for any $x\in W_{i}$ and any class $W_{i}$. We conclude that $A$ is a weak point attractor. ∎
Step (II) has already been stated in [EOMUW24, Theorem 9], for which we present a simplified proof.
Proposition 3.4 (II).
$A$ is a forward point attractor.
Proof.
If $\varphi^{N}_{\omega}(x)\in A(\theta_{N}\omega)$, then also $\varphi^{n}_{\omega}(x)\in A(\theta_{n}\omega)$ for any $n\geq N$. Using continuity from below, we find
$\mathds{P}\big{(}\exists n\in\mathds{N}:\varphi^{n}_{\omega}(x)\in A(\theta_{n}\omega)\big{)}$ $=\lim_{n\to\infty}\mathds{P}\big{(}\varphi^{n}_{\omega}(x)\in A(\theta_{n}\omega)\big{)}=1.$
The second equality uses that $A$ is a weak point attractor. This shows that $A$ is a forward point attractor. ∎
Proposition 3.5 (III).
$A$ is a pullback point attractor.
Proof.
By the correspondence theorem, the pullback-limit of $\pi$ converges to $\mu_{\omega}$ for $\mathds{P}$-a.e. $\omega\in\Omega$, i.e.
$\lim_{n\to\infty}[(\varphi^{n}_{\theta_{-{n}}\omega})_{*}\pi](B)=\mu_{\omega}(B),$ (18)
for any $B\subset\mathds{X}$. By equality (15), we have $\mu_{\omega}(A(\omega))=1$ for $\mathds{P}$-a.e. $\omega$. Hence, equation (18) yields
$\lim_{n\to\infty}[(\varphi^{n}_{\theta_{-{n}}\omega})_{*}\pi](A(\omega))=1\,,\quad\text{ for $\mathds{P}$-a.e.~{}$\omega$.}$ (19)
Assume now for contradiction that $A$ is not a pullback point attractor. Then there are an $x\in\mathds{X}$ and a set $\Omega^{x}\subset\Omega$ with $\mathbb{P}(\Omega^{x})>0$ such that for every $\omega\in\Omega^{x}$ we have an increasing sequence $n_{k}\to\infty$ such that for each $k\in\mathds{N}$
$\varphi^{n_{k}}_{\theta_{-{n_{k}}}\omega}(x)\notin A(\omega).$
This implies that $[(\varphi^{n_{k}}_{\theta_{-{n_{k}}}\omega})_{*}\pi](A(\omega))\leq 1-\pi(x)<1$ for any $\omega\in\Omega^{x}$ and $k\in\mathbb{N}$. This is a contradiction to (19) since $\Omega^{x}$ has positive measure. ∎
Proposition 3.6 (IV).
In countable state spaces pullback point attraction is equivalent to pullback attraction for finite deterministic sets and forward point attraction is equivalent to forward attraction of compact random sets.
Proof.
Let $A$ be a pullback point attractor. We show that $A$ is a pullback attractor for finite deterministic sets.
Let $B=\{x_{1},\ldots,x_{m}\}\subset\mathds{X}$ be a finite set. Since $A$ is a pullback point attractor, for $\mathds{P}$-a.e. $\omega\in\Omega$ there are numbers $N_{1}(\omega),\ldots,N_{m}(\omega)$ such that for all $n\geq N_{i}(\omega)$, we find $\varphi^{n}_{\theta_{-n}\omega}(x_{i})\subset A(\omega)$. Let $N(\omega)=\max\{N_{1}(\omega),\ldots,N_{m}(\omega)\}$, implying $\varphi^{n}_{\theta_{-n}\omega}(B)\subset A(\omega)$ for all $n\geq N(\omega)$. We conclude that
$\mathds{P}(\exists N\in\mathds{N}:\forall n\geq N:\varphi^{n}_{\theta_{-n}\omega}(B)\subset A(\omega))=1.$
This shows that $A$ is a pullback attractor for finite deterministic sets.
Let now $A$ be a forward point attractor. By an analogous argument as above, $A$ is a forward attractor for finite deterministic sets. Let $B$ be a compact random set and denote the set of finite subsets of $\mathds{X}$ by $\mathcal{P}_{c}(\mathds{X})$, which is countable. We compute
$\mathds{P}(\exists N\in\mathds{N}:\forall n\geq N:\varphi^{n}_{\omega}(B(\omega))\subset A(\theta_{n}\omega))$
$=\sum_{K\in\mathcal{P}_{c}(\mathds{X})}$ $\mathds{P}(\exists N\in\mathds{N}:\forall n\geq N:\varphi^{n}_{\omega}(K)\subset A(\theta_{n}\omega)\land B(\omega)=K)$
$=\sum_{K\in\mathcal{P}_{c}(\mathds{X})}$ $\mathds{P}(B(\omega)=K)=1,$
where we used that $A$ is a forward attractor for finite deterministic sets. This shows that $A$ is a forward attractor for compact random sets.
Proposition 3.7 (V).
There is a constant $\kappa\in\mathds{N}$ such that $\left|A(\omega)\right|=\kappa$ for $\mathds{P}$-a.e. $\omega\in\Omega$. The probability that a state $x\in\mathds{X}$ lies in the attractor is given by
$\mathds{P}(x\in A(\omega))=\kappa\pi(x).$
Proof.
By (14), there is a constant $\kappa\in\mathds{N}$ such that the sample measures of the unique invariant Markov measure $\mu$ are given by
$\mu_{\omega}=\frac{1}{\kappa}\sum_{a\in A(\omega)}\delta_{a}.$
Hence, the attractor $A(\omega)$ consists of exactly $\kappa$ points for $\mathds{P}$-a.e. $\omega\in\Omega$. We compute
$\mathds{E}[\mu_{\omega}(x)]=\frac{1}{\kappa}\mathds{P}(x\in A(\omega)).$
By the correspondence theorem, we have $\mathds{E}[\mu_{\omega}]=\pi$, and the statement follows. ∎
3.1 Examples
We provide an example of a Markov chain on a finite state space that corresponds with two different random dynamical systems with attractors of different cardinality, respectively.
Example 1.
(RDS representations with attractors of different sizes)
Let $\mathds{X}=\{a,b,c,d\}$. Consider the Markov chain with transition probabilities $P$ as indicated by the following diagram.
$a$$b$$c$$d$$\frac{1}{2}$$\frac{1}{2}$$\frac{1}{2}$$\frac{1}{2}$$\frac{1}{2}$$\frac{1}{2}$$\frac{1}{2}$$\frac{1}{2}$ (20)
We present two possible RDS representations of this Markov chain.
We first consider the independent RDS $(\theta,\varphi)$, cf. Section 2.4. In each timestep any point may either take a step clockwise, or remain in its position. Since in the independent RDS all points move independently from one another, there are 16 possible maps $\varphi_{\omega}:\mathds{X}\to\mathds{X}$ that are taken with positive probability. Since all transition probabilities are $\frac{1}{2}$, each of these 16 possible maps occurs with probability $\frac{1}{16}$. It is not hard to verify that with probability $1$ all points eventually coalesce. Hence, the unique attractor $A(\omega)$ consists $\mathds{P}$-a.s. of a single point.
The second RDS representation we present is constructed over a shift space, like described in Section 2.4. The map $\varphi_{\omega}:\mathds{X}\to\mathds{X}$ admits one of the following two configurations, each with probability $\frac{1}{2}$.
$a$$b$$c$$d$$f_{1}$$a$$b$$c$$d$$f_{2}$ (21)
The RDS $(\theta,\varphi)$ over the shift space constructed by $\mathds{P}_{0}(f=f_{1})=\mathds{P}_{0}(f=f_{2})=\frac{1}{2}$ represents the Markov chain in (20). Note that points diagonal to one another never coalesce. After one time step, all points lie on one of the two diagonals $\{a,c\}$ or $\{b,d\}$. Given $\omega=(\ldots,\omega_{-1},\omega_{0},\omega_{1},\ldots)\in\Omega$ with $\omega_{i}\in\{f_{1},f_{2}\}$, define
$A(\omega)=\begin{cases}\{a,c\},\quad&\textup{if }\omega_{-1}=f_{2},\\ \{b,d\},\quad&\textup{if }\omega_{-1}=f_{1}.\end{cases}$
By construction, $A$ is a compact random invariant set. Additionally, $A$ absorbs any initial point of $\mathds{X}$ within one time step. Hence, $A$ is the unique attractor which is $\mathds{P}$-a.s. of size 2. $\Diamond$
The proofs of Proposition 3.4 and Proposition 3.6 do not use the positive recurrence or irreducibility of the Markov chain. Hence, the notions of forward attraction, weak attraction and their respective point attractions are equivalent purely by the discrete-time, discrete-space setting. The following example shows that, without the assumption of a unique stationary distribution with full support, these types of attractors are in general not pullback attractors.
Example 2.
(Forward but not pullback attractor) Consider a random walk on $\mathds{X}=\mathds{N}_{0}$ that is absorbing in 00. The transition probabilities are given by $P(x,x+1)=P(x,x-1)=\frac{1}{2}$ for $x\in\mathds{N}$ and $P(0,0)=1$.
We consider the RDS in which in each time step either all non-zero points move up or all non-zero points move down. We construct this RDS over a shift-space, like described in Section 2.4. The map $\varphi_{\omega}$ admits one of the following two configurations, each with probability $\frac{1}{2}$:
$f_{1}(x)=\begin{cases}x+1,\quad&x>0,\\ 0,&x=0,\end{cases}\qquad f_{-1}(x)=\begin{cases}x-1,\quad&x>0,\\ 0,&x=0.\end{cases}$
We write elements $\omega\in\Omega$ as $\omega=(\ldots,\omega_{-1},\omega_{0},\omega_{1},\ldots)$, where $\omega_{n}\in\{f_{-1},f_{1}\}$. It is well-known that the random walk $X_{n}:=\varphi^{n}_{\omega}(x)$ eventually ends up in the absorbing state $\{0\}$ for $\mathds{P}$-a.e. $\omega\in\Omega$, e.g. [MT12, Section 8.1]. Hence, the invariant compact random set given by $A(\omega)=\{0\}$ is a forward point attractor. By Proposition 3.6, $A$ is even a forward attractor for compact random sets. However, we will see that $A$ is not a pullback point attractor.
For $\omega\in\Omega$ and $n\in\mathds{N}$ consider the set
$C_{n}(\omega):=\{x\in\mathds{X}\mid\varphi_{\theta_{-n}\omega}^{n}(x)=0\}.$
Note that the RDS is monotone in the sense that if $x\leq y$, then $\mathds{P}$-a.s. also $\varphi_{\omega}^{n}(x)\leq\varphi_{\omega}^{n}(y)$. Hence, the sets $C_{n}(\omega)$ are of the form $\{0,\ldots,m_{n}(\omega)\}$. We claim that $(m_{n}(\omega))_{n\in\mathds{N}}$ is itself a random walk on $\mathds{N}_{0}$. By the cocycle property we have
$x\in C_{n+1}(\omega)\quad\iff\quad\varphi_{\theta_{-(n+1)}\omega}^{1}(x)\in C_{n}(\omega).$
Hence, we have a recursive formula for $(m_{n}(\omega))$,
$m_{n+1}(\omega)=\begin{cases}m_{n}(\omega)+1,\quad&\omega_{-(n+1)}=f_{-1},\\ m_{n}(\omega)-1,\quad&\omega_{-(n+1)}=f_{1}\>\land\>m_{n}(\omega)>0,\\ 0,&\omega_{-(n+1)}=f_{1}\>\land\>m_{n}(\omega)=0.\end{cases}$
A visualization of a few trajectories of this RDS and the evolution of $m_{n}(\omega)$ is depicted in Figure 1. Since the elements $\omega_{n}$ are independent from one another, $(m_{n}(\omega))_{n\in\mathds{N}}$ is a Markov chain on $\mathds{N}_{0}$ with transition kernel $Q(m,m+1)=Q(m,m-1)=\frac{1}{2}$ for $m\in\mathds{N}$ and $Q(0,0)=Q(0,1)=\frac{1}{2}$. This Markov chain is a classic random walk on $\mathds{N}_{0}$ that is known to be recurrent, e.g. [MT12, Section 8.1]. Therefore, for $\mathds{P}$-a.e. $\omega\in\Omega$, there is an increasing sequence $n_{k}$ such that $m_{n_{k}}(\omega)=0$. By definition of $m_{n}(\omega)$, this implies $\varphi_{\theta_{-n_{k}}\omega}^{n_{k}}(x)\neq 0$ for all $x\in\mathds{N}$. We conclude that $A$ is not a pullback point attractor. $\Diamond$
[…]
Figure 1: A visualization of trajectories of the RDS from Example 2. The $x$-axis denotes time; the $y$-axis is the state space $\mathds{X}=\mathds{N}_{0}$. An arrow from $(\theta_{-n}\omega,x)$ to $(\theta_{-n+1}\omega,y)$ indicates that $\varphi_{\theta_{-n}\omega}^{1}x=y$. The thick line shows the value of $m_{n}(\omega)$. The depicted graph on the set of nodes ×$\mathds{Z}\times\mathds{X}$ has been called Doeblin graph in the literature [BHMM19].
4 Insulated and synchronizing sets
In the following, we analyze the synchronizing behavior for sets of initial points, also characterizing the structure of the unique attractor $A$, i.e. the number of elements of $A(\omega)$ and their configuration. We stress again that the RDS representation of a Markov chain is not unique. While the existence of the unique random attractor $A$ does not depend on the choice of the RDS representation (cf. Theorem 3.1), the structure of the attractor $A$ certainly does (cf. Example 1).
4.1 Insulated sets
Let $(\theta,\varphi)$ be a fixed RDS representation of an irreducible and positive recurrent Markov chain with transition probabilities $P$ and stationary distribution $\pi$. We introduce the following relation on $\mathds{X}$.
Definition 4.1.
We call two points $x,y\in\mathds{X}$ insulated, written $x\parallel y$, if they do not $\omega$-synchronize for $\mathds{P}$-a.e. $\omega\in\Omega$, i.e.
$\forall n\in\mathds{N}:\mathds{P}(\varphi^{n}_{\omega}(x)=\varphi^{n}_{\omega}(y))=0.$
An insulated set is a set $K\subset\mathds{X}$ such that $x\parallel y$ for any $x\neq y\in K$. We denote the size of the largest insulated set by $\hat{\kappa}\in\mathds{N}\cup\{\infty\}$. We call an insulated set $K\subset\mathds{X}$ with $|K|=\hat{\kappa}$ a maximum insulated set.
Remark 4.2.
  1. (i)
    There might be insulated sets $K\subset\mathds{X}$ with $|K|<\hat{\kappa}$ such that adding any point $x$ to $K$ makes $K\cup\{x\}$ not insulated. Such a set might be called a maximal insulated set, even though it is not a maximum insulated set. Maximal insulated sets will play no role in the rest of this work. Still, this distinction is important to avoid confusion in the terminology.
  2. (ii)
    In the previous section, we showed that the RDS $(\theta,\varphi)$ has a unique attractor $A$ of size $\left|A(\omega)\right|=\kappa$ for $\mathds{P}$-a.e. $\omega\in\Omega$. Theorem 4.5 below shows that, in fact, $\hat{\kappa}=\kappa$. However, until this theorem is proven, we denote the size of the maximum anticlique by $\hat{\kappa}$ to distinguish it from the $\mathds{P}$-a.s. size of the attractor.
For an example of insulated sets of an RDS, see Example 1, in particular (21). The next result states that insulated pairs of points remain so under the random dynamics.
Proposition 4.3.
Let $x,y\in\mathbb{X}$ such that $x\parallel y$. Then $\varphi_{\omega}(x)\parallel\varphi_{\omega}(y)$ for $\mathds{P}$-a.e. $\omega\in\Omega$.
Proof.
Let us assume for contradiction that $\varphi_{\omega}(x)\nparallel\varphi_{\omega}(y)$ with positive probability. Since the set of pairs $\{(x^{\prime},y^{\prime})\in\mathbb{X}\times\mathbb{X}\,:\,x^{\prime}\nparallel y^{\prime}\}$ is countable, there are $x^{\prime},y^{\prime}\in\mathbb{X}$ with $x^{\prime}\nparallel y^{\prime}$ such that the set
$\Omega^{(x^{\prime},y^{\prime})}:=\{\omega\in\Omega\>|\>\varphi_{\omega}(x,y)=(x^{\prime},y^{\prime})\}$
has positive probability. Since $x^{\prime}\nparallel y^{\prime}$, there is an $n\in\mathds{N}$ such that the set
$\Omega^{n}:=\{\tilde{\omega}\in\Omega\>|\>\varphi^{n}_{\tilde{\omega}}(x^{\prime})=\varphi^{n}_{\tilde{\omega}}(y^{\prime})\}$
has positive probability. Note that $\Omega^{(x^{\prime},y^{\prime})}$ is $\mathcal{F}_{0}^{1}$-measurable while $\Omega^{n}$ is $\mathcal{F}_{0}^{\infty}$-measurable. In particular $\theta_{-1}\Omega^{n}$ is $\mathcal{F}_{1}^{\infty}$-measurable, and thereby independent of $\Omega^{(x^{\prime},y^{\prime})}$. Hence,
$\mathds{P}(\Omega^{(x^{\prime},y^{\prime})}\cap\theta_{-1}\Omega^{n})>0.$
Let $\omega\in\Omega^{(x^{\prime},y^{\prime})}\cap\theta_{-1}\Omega^{n}$ such that $\tilde{\omega}:=\theta\omega\in\Omega^{n}$. For such $\omega$ we find
$\varphi_{\omega}^{n+1}(x)=\varphi_{\theta\omega}^{n}\circ\varphi_{\omega}(x)=\varphi_{\tilde{\omega}}^{n}(x^{\prime})=\varphi_{\tilde{\omega}}^{n}(y^{\prime})=\varphi_{\theta\omega}^{n}\circ\varphi_{\omega}(y)=\varphi_{\omega}^{n+1}(y).$
This shows that with positive probability there is an $n\in\mathds{N}$ such that $\varphi_{\omega}^{n}(x)=\varphi_{\omega}^{n}(y)$. This contradicts $x\parallel y$. Hence $\varphi_{\omega}(x)\parallel\varphi_{\omega}(y)$ for $\mathds{P}$-a.e. $\omega\in\Omega$. ∎
A direct consequence of this lemma is the following important observation.
Corollary 4.4.
Let $K\subset\mathds{X}$ be an insulated set. For any $n\in\mathds{N}$ and $\mathds{P}$-a.e. $\omega\in\Omega$ the set $\varphi^{n}_{\omega}(K)$ is also an insulated set, with $|\varphi^{n}_{\omega}(K)|=|K|$.
Proof.
By Proposition 4.3, for each pair $x\parallel y$, there is a measurable set $\Omega^{(x,y)}\subset\Omega$ of full probability such that $\varphi_{\omega}(x)\parallel\varphi_{\omega}(y)$ for all $\omega\in\Omega^{(x,y)}$. Define the set
$\overline{\Omega}:=\bigcap_{x\parallel y}\bigcap_{n\in\mathds{Z}}\theta_{n}\Omega^{(x,y)}.$ (22)
As a countable intersection of measurable sets of full probability, $\overline{\Omega}$ has full probability. For $\omega\in\overline{\Omega}$, we find $\varphi_{\omega}^{n}(x)\parallel\varphi_{\omega}^{n}(y)$ for any $n\in\mathds{N}$ and $x\parallel y$. In particular, for any insulated set $K\subset\mathds{X}$ and $n\in\mathds{N}$, we get that $\varphi_{\omega}^{n}(K)$ is an insulated set of size $|K|$. This completes the proof. ∎
The following theorem now establishes the random attractor as a maximum insulated set.
Theorem 4.5.
For $\mathds{P}$-a.e. $\omega\in\Omega$, the attractor $A(\omega)$ is a maximum insulated set.
Proof.
By definition of a weak attractor, we have for any finite set $K$ that
$\lim_{n\to\infty}\mathds{P}(\varphi^{n}_{\omega}(K)\subset A(\theta_{n}\omega))=1.$
If $K$ is a maximum insulated set, then, by Corollary 4.4, for $\mathds{P}$-a.e. $\omega\in\Omega$ and any $n\in\mathds{N}$ the set $\varphi^{n}_{\omega}(K)$ is a maximum insulated set as well. Hence, using the measure-preserving property of $\theta$, the set $A(\omega)$ contains a maximum insulated set $\mathds{P}$-a.s. and, therefore, $|A(\omega)|\geq\hat{\kappa}$.
In the proof of Proposition 3.3, we already showed that $A(\omega)$ is $\mathds{P}$-a.s. of constant finite size $\kappa$. Assume now for contradiction that $\kappa>\hat{\kappa}$ for $\mathds{P}$-a.e. $\omega\in\Omega$. therefore, for each such $\omega$, the set $A(\omega)$ contains at least two points that are not insulated. Since the set of subsets of $\mathds{X}$ of size $\kappa$ is countable, there are points $x\nparallel y$ such that the measurable set
$\Omega^{(x,y)}:=\{\omega\in\Omega\>|\>x,y\in A(\omega)\}$
has positive probability. Since $x\nparallel y$, there is an $n\in\mathds{N}$ such that $\varphi_{\omega}^{n}(x)=\varphi_{\omega}^{n}(y)$ with positive probability, and we can define the measurable set
$\Omega^{n}=\{\omega\in\Omega\>|\>\varphi^{n}_{\omega}(x)=\varphi^{n}_{\omega}(y)\}$
such that $\mathbb{P}\left(\Omega^{n}\right)>0$. Observe that $\Omega^{(x,y)}$ is $\mathcal{F}_{-\infty}^{0}$-measurable, while $\Omega^{n}$ is $\mathcal{F}_{0}^{n}$-measurable. Hence, the two events are independent and their intersection has positive probability. For $\omega\in\Omega^{(x,y)}\cap\Omega^{n}$, the points $x,y$, which both lie in $A(\omega)$, coalesce after $n$ steps. Hence, we obtain
$|A(\theta_{n}\omega)|=\left|\varphi^{n}_{\omega}(A(\omega)\right|<\kappa=|A(\omega)|,$
which is a contradiction. Hence, $A(\omega)$ is $\mathds{P}$-a.s. a maximum insulated set with $|A(\omega)|=\hat{\kappa}$. ∎
Remark 4.6.
  1. (i)
    Theorem 4.5 shows that the $\mathds{P}$-a.s. size $\kappa$ of the attractor equals $\hat{\kappa}$, the size of a maximum insulated set. Hence, we use $\kappa$ from now on to denote this quantity. Example 1 illustrates that the value of $\kappa$ may differ for different RDS representations of the same Markov chain. The independent RDS always satisfies $\kappa=p$, where $p$ is the period of the Markov chain.
  2. (ii)
    The constant $\kappa$ determines whether a given RDS is synchronizing or not. If $\kappa>1$, the RDS cannot be synchronizing since points in a maximum insulated set $\mathds{P}$-a.s. never synchronize. If $\kappa=1$, then any two points $x,y\in\mathds{X}$ eventually hit the attractor $A(\omega)$ which consists of a single point. At that time $x$ and $y$ have synchronized. Hence, an RDS is synchronizing if and only if $\kappa=1$.
4.2 Synchronizing sets
The previous subsection addressed subsets of points in $\mathds{X}$ that do not $\omega$-synchronize for $\mathds{P}$-a.e. $\omega\in\Omega$. In this subsection, we characterize the sets of points that do $\omega$-synchronize. Recall that two points $x,y\in\mathds{X}$ are said to $\omega$-synchronize if there is an $n\in\mathds{N}$ such that $\varphi_{\omega}^{n}(x)=\varphi_{\omega}^{n}(y)$.
Given $\omega\in\Omega$, define the equivalence relation
[…]
We denote the equivalence classes by $[x]_{\omega}$. We call these $\omega$-dependent equivalence classes synchronizing classes instead. Note that $x\parallel y$ is equivalent to $\mathds{P}(x\sim_{\omega}y)=0$. In general, we can have $10<\mathds{P}(x\sim_{\omega}y)<$ 1, cf. Example 1.
Proposition 4.7.
For $\mathds{P}$-a.e. $\omega\in\Omega$, there are exactly $\kappa$ synchronizing sets. Each synchronizing set $[x]_{\omega}$ satisfies $\pi([x]_{\omega})=\frac{1}{\kappa}$.
Proof.
Since the attractor $A$ is a forward point attractor, each $x\in\mathds{X}$ $\omega$-synchronizes with one of the elements of $A(\omega)$, i.e. $x\in[a]_{\omega}$ for some $a\in A(\omega)$, for $\mathds{P}$-a.e. $\omega\in\Omega$. Since $A(\omega)$ consists of $\kappa$ elements for $\mathds{P}$-a.e. $\omega\in\Omega$, there are at most $\kappa$ synchronizing sets. On the other hand, insulated points $\mathds{P}$-a.s. lie in different synchronizing sets. Hence, for $\mathds{P}$-a.e. $\omega\in\Omega$ there are exactly $\kappa$ synchronizing sets.
To prove the second statement, fix $\tilde{\omega}\in\Omega$ such that there are exactly $\kappa$ synchronizing sets and each maximum insulated set has exactly one element inside each synchronizing set. This is the case for $\mathds{P}$-a.e. $\tilde{\omega}$. Let $[x]_{\tilde{\omega}}$ be one of the synchronizing sets. By the correspondence theorem, we find $\pi([x]_{\tilde{\omega}})=\mathds{E}[\mu_{\omega}([x]_{\tilde{\omega}})]$, where the expectation is taken over $\omega$. The sample measures $\mu_{\omega}$ are a uniform distribution over the sets $A(\tilde{\omega})$, cf. (14), which yields
$\pi([x]_{\tilde{\omega}})=\mathds{E}[\mu_{\omega}([x]_{\tilde{\omega}})]=\int_{\Omega}\frac{1}{\kappa}\sum_{a\in A(\omega)}\delta_{a}([x]_{\tilde{\omega}})\;\mathds{P}(d\omega).$
For $\mathds{P}$-a.e. $\omega\in\Omega$, the set $A(\omega)$ is a maximum insulated set. Hence, exactly one of the summands is $1$ while the others are 00. We conclude $\pi([x]_{\tilde{\omega}})=\frac{1}{\kappa}$. This holds for $\mathds{P}$-a.e. $\tilde{\omega}\in\Omega$, which completes the proof.
Remark 4.8.
Proposition 4.7 imposes significant restrictions on the value of $\kappa$. If it is not possible to divide $\mathds{X}$ into $\ell$ sets of equal $\pi$-mass, then $\kappa\neq\ell$. This is relevant since the distribution of $\pi$ depends solely on the associated Markov chain and not on the RDS representation.
4.3 Dynamics on maximum insulated sets
Theorem 4.5 shows that the attractor $A(\omega)$ is a maximum insulated set for almost all $\omega\in\Omega$. Hence, to understand the structure of the random attractor in more detail, we study the random dynamics restricted to the set of maximum insulated sets. The results of this section will be used in the proofs of Section 5.4.
Let $\mathcal{I}\subset\mathcal{P}(\mathds{X})$ be the set of maximum insulated subsets. Let $\mathcal{A}\subset\mathcal{I}$ be the subset of all maximum insulated sets which the attractor $A$ admits with positive probability, i.e.
$\mathcal{A}:=\{K\in\mathcal{I}\>|\>\mathds{P}(A(\omega)=K)>0\}.$ (23)
For $\mathds{P}$-a.e. $\omega\in\Omega$, we have, by the Markov property,
$K\in\mathcal{I}\>\Rightarrow\>\varphi_{\omega}(K)\in\mathcal{I},\qquad K\in\mathcal{A}\>\Rightarrow\>\varphi_{\omega}(K)\in\mathcal{A}.$
Similarly to (22), we can define a measurable, $\theta$-invariant set $\overline{\Omega}\subset\Omega$ of full probability such that
$K\in\mathcal{I}\Rightarrow\varphi^{n}_{\omega}(K)\in\mathcal{I},\qquad K\in\mathcal{A}\Rightarrow\varphi^{n}_{\omega}(K)\in\mathcal{A},\qquad A(\omega)\in\mathcal{A},$ (24)
for all $\omega\in\overline{\Omega}$ and $n\in\mathds{N}$. Hence, the Markov RDS $(\theta,\varphi)$ can be interpreted as a Markov RDS over $\overline{\Omega}$ on the state space $\mathcal{I}$, with absorbing, i.e. forward-invariant, set $\mathcal{A}$. This RDS defines a Markov chain on the state space $\mathcal{I}$ with transition probabilities
$S(K,L):=\mathds{P}(\varphi_{\omega}(K)=L),$ (25)
for $K,L\in\mathcal{I}$. Since $\mathcal{A}$ is absorbing, this Markov chain is not irreducible unless $\mathcal{A}=\mathcal{I}$. Recall that an irreducibility class is a subset $\mathcal{C}\subset\mathcal{I}$ such that $S(K,\mathcal{C})=1$ for each $K\in\mathcal{C}$ and the Markov chain restricted to $\mathcal{C}$ is irreducible. Let $\tau_{\mathcal{A}}$ be the hitting time of the set $\mathcal{A}$. We denote the probability measure corresponding to this Markov chain, started in $K\in\mathcal{I}$, by $\mathds{P}^{S}_{K}$.
Theorem 4.9.
  1. (i)
    The set $\mathcal{A}$ is the unique irreducibility class of the Markov chain defined via (25) on the state space $\mathcal{I}$.
  2. (ii)
    Each $K\in\mathcal{I}$ eventually enters the set $\mathcal{A}$, i.e. $\mathds{P}_{K}^{S}(\tau_{\mathcal{A}}<\infty)=1$.
  3. (iii)
    Restricted to $\mathcal{A}$, the Markov chain is positive recurrent with unique stationary distribution
    $\varrho(K):=\mathds{P}(A(\omega)=K),\qquad\forall K\in\mathcal{A}.$
Proof.
(i) The fact that $S(K,\mathcal{A})=1$ for every $K\in\mathcal{A}$ is clear from above. Furthermore, for each $K,L\in\mathcal{A}$, define the events
$E:=\{\omega\in\overline{\Omega}\>|\>A(\omega)=K\},\qquad F:=\{\omega\in\overline{\Omega}\>|\>A(\omega)=L\}.$
By definition, both of these events have positive probability. The Poincaré recurrence theorem states that
$\mathds{P}(\theta_{n}\omega\in F\>\text{for infinitely many }n\in\mathds{N})=1.$
Hence, there is an $n\in\mathds{N}$ such that $\mathds{P}(\omega\in E\land\theta_{n}\omega\in F)>0$, which by the invariance of $A$, it can be rewritten as
$\mathds{P}(\omega\in E\land\theta_{n}\omega\in F)$ $=\mathds{P}(A(\omega)=K\land\varphi_{\omega}^{n}(A(\omega))=L)$
$=\mathds{P}(\varphi_{\omega}^{n}(A(\omega))=L\>|\>A(\omega)=K)\>\mathds{P}(A(\omega)=K)$
$=S^{n}(K,L)\>\mathds{P}(A(\omega)=K).$
Since the left hand side is positive, this shows $S^{n}(K,L)>0$ and, hence, $\mathcal{A}$ is an irreducibility class.
Assume there was another irreducibility class $\mathcal{C}\subset\mathcal{I}\setminus\mathcal{A}$. Let $K\in\mathcal{C}$. Since $S(K,\mathcal{C})=1$, we have $S^{n}(K,\mathcal{A})=0$ for any $n\in\mathds{N}$. In particular, we find
$\mathds{P}(\varphi_{\omega}^{n}(K)=A(\theta_{n}\omega))\leq\mathds{P}(\varphi_{\omega}^{n}(K)\in\mathcal{A})=S^{n}(K,\mathcal{A})=0.$
This is a contradiction to $A$ being a weak attractor, cf. (11). Hence, $\mathcal{A}$ is the unique irreducibility class of the Markov chain on $\mathcal{I}$.
(ii) Since $A(\omega)\in\mathcal{A}$ for $\omega\in\overline{\Omega}$ we find
$\mathds{P}^{S}_{K}(\tau_{\mathcal{A}}\leq n)\geq\mathds{P}(\varphi_{\omega}^{n}(K)\in A(\theta_{n}\omega)).$
Since $A$ is a weak forward attractor for finite deterministic sets, the right-hand side approaches $1$ as $n\to\infty$. Hence, $\mathds{P}_{K}^{S}(\tau_{\mathcal{A}}<\infty)=1$.
(iii) To verify that $\varrho$ is a stationary distribution we compute
$(\varrho S)(L)$ $=\sum_{K\in\mathcal{A}}\varrho(K)S(K,L)$
$=\sum_{K\in\mathcal{A}}\mathds{P}(A(\omega)=K)\>\mathds{P}(\varphi_{\omega}(K)=L)$
$=\sum_{K\in\mathcal{A}}\mathds{P}(A(\omega)=K)\>\mathds{P}(A(\theta\omega)=L\>|\>A(\omega)=K)$
$=\sum_{K\in\mathcal{A}}\mathds{P}(A(\theta\omega)=L\land A(\omega)=K)$
$=\mathds{P}(A(\theta\omega)=L)=\varrho(L).$
In the last line, we used the the measure-preserving property of $\theta$. This completes the proof. ∎
5 Expected hitting times
Let $(\theta,\varphi)$ be a fixed RDS representation of an irreducible and positive recurrent Markov chain with transition probabilities $P$ and stationary distribution $\pi$. Theorem 3.1 states that there is a unique attractor $A$. Consider the random hitting time of the attractor for an initial condition $x\in\mathds{X}$
$T_{A}(\omega,x):=\min\{n\in\mathds{N}\>|\>\varphi^{n}_{\omega}(x)\in A(\theta_{n}\omega)\}.$ (26)
Since $A$ is a forward attractor, the random time $T_{A}(\omega,x)$ is $\mathds{P}$-a.s. finite. In Section 5.1, we provide necessary and sufficient conditions for the expectation $\mathds{E}[T_{A}(\omega,x)]$ to be finite for all $x\in\mathds{X}$, cf. Theorem 5.2.
In Section 5.2, we restrict ourselves to synchronizing RDS, i.e. $\kappa=1$. In that case the synchronization time
$T_{\Delta}(\omega,x,y):=\min\{n\in\mathds{N}\mid\varphi_{\omega}^{n}(x)=\varphi_{\omega}^{n}(y)\},$ (27)
is $\mathds{P}$-a.s. finite for any $x,y\in\mathds{X}$. Theorem 5.4 shows that the expectation $\mathds{E}[T_{\Delta}(\omega,x,y)]$ is always finite for all $x,y\in\mathds{X}$, even when the associated Markov chain is not ergodic of degree 2.
5.1 Time until attraction
In this section, we consider the time $T_{A}(\omega,x)$ it takes for a fixed point $x\in\mathds{X}$ to hit the random attractor $A$, cf. (26). Since $A$ is a forward point attractor, this time is $\mathds{P}$-a.s. finite. However, we will see below, cf. Example 3, that $T_{A}(\omega,x)$ does not necessarily have finite expectation. The main result of this section, Theorem 5.2 below, states that $T_{A}(\omega,x)$ is of finite mean for all $x\in\mathds{X}$ if and only if the associated irreducible, positive recurrent Markov chain is ergodic of degree 2, i.e. for some $x\in\mathds{X}$ we have
$\mathds{E}_{x}[\tau^{2}_{x}]<\infty,$
where $\tau_{x}$ is the first hitting time of $x$ as given in (2). The following proposition lists equivalent characterizations of ergodicity of degree 2 which we will use in the following. For alternative characterizations of ergodicity of degree 2 and higher degrees, we refer to [Mao03].
Proposition 5.1.
Consider an irreducible, positive recurrent Markov chain with stationary distribution $\pi$. The following are equivalent:
  1. (i)
    The Markov chain is ergodic of degree 2.
  2. (ii)
    $\mathds{E}_{x}[\tau_{y}^{2}]<\infty$, for all $x,y\in\mathds{X}$.
  3. (iii)
    $\mathds{E}_{\pi}[\tau_{y}]=\sum_{x\in\mathds{X}}\pi(x)\mathds{E}_{x}[\tau_{y}]<\infty$, for all $y\in\mathds{X}$.
Proof.
The equivalence of $(i)$ and $(ii)$ is shown in [Mao03, Remark 3.1]. We show the equivalence of $(i)$ and $(iii)$. Due to [BH22, Proposition 2.10], we have the identity
$\mathds{E}_{\pi}[\tau_{y}]=\frac{1}{2}\pi(y)(\mathds{E}_{y}[\tau^{2}_{y}]+\mathds{E}_{y}[\tau_{y}]).$
Since $\mathds{E}_{y}[\tau_{y}]<\infty$ for positive recurrent Markov chains, we find $\mathds{E}_{\pi}[\tau_{y}]<\infty$ if and only if $E_{y}[\tau_{y}^{2}]<\infty$. ∎
We are now ready to state the main theorem of the section.
Theorem 5.2.
The expected hitting time of the attractor $\mathds{E}[T_{A}(\omega,x)]$ is finite for all $x\in\mathds{X}$ if and only if the associated Markov chain is ergodic of degree 2. In that case, we even have
$\sum_{x\in\mathds{X}}\pi(x)\mathds{E}[T_{A}(\omega,x)]<\infty.$ (28)
For a proof of this theorem, see Section 5.4.
Remark 5.3.
Theorem 5.2 shows that the finiteness of the expected time until attraction for all $x\in\mathds{X}$ does not depend on the RDS representation, but only on the associated Markov chain. However, even for Markov chains which are ergodic of degree 2, the time until attraction may still be arbitrarily high depending on the RDS representation. In other words, there is no upper bound on the quantity in (28) which does not depend on the RDS representations. This is demonstrated in Example 4.
We emphasize that Theorem 5.2 states that the associated Markov chain is ergodic of degree 2 if and only if $\mathds{E}[T_{A}(\omega,x)]$ is finite for all $x\in\mathds{X}$. So far, we were unable to exclude the case that in an RDS representing a Markov chain that is not ergodic of degree 2 there is one point $x\in\mathds{X}$ with $\mathds{E}[T_{A}(\omega,x)]<\infty$ and another point $y\in\mathds{X}$ with $\mathds{E}[T_{A}(\omega,y)]=\infty$. We propose the following conjecture.
Conjecture 1.
The expected time until attraction $\mathds{E}[T_{A}(\omega,x)]$ is finite for one $x\in\mathds{X}$ if and only it is finite for all $x\in\mathds{X}$.
In Section 5.2, we prove the conjecture in the special case of synchronizing RDS.
We provide some intuition for why ergodicity of degree 2 is the necessary and sufficient condition for obtaining attraction in a time of finite mean: The attractor $A$ is distributed according to $\mathds{P}(x\in A(\omega))=\kappa\pi(x)$, cf. (13). In particular, for $\kappa=1$, the random point $A(\omega)$ is distributed according to $\pi$. If a state $y$ hits the attractor in a time of finite mean, then, since the trajectory of $y$ returns to $y$ in a time of finite mean, the time it takes for the attractor to reach the state $y$ is of finite mean as well. This is possible only if the underlying Markov chain is ergodic of degree 2, as can be seen from the characterization in Proposition 5.1 (iii).
5.2 Finite expected synchronization time
In this section, we consider synchronizing RDS, i.e. the case $\kappa=1$. In particular, any two initial states $\omega$-synchronize, for $\mathds{P}$-a.e. $\omega\in\Omega$. Hence, the synchronization time $T_{\Delta}(\omega,x,y)$, cf. (12), is $\mathds{P}$-a.s. finite for any $x,y\in\mathds{X}$. We show that the expectation of this random time is finite as well. Theorem 5.2 from the previous section states that if the associated Markov chain is ergodic of degree 2, any starting point $x\in\mathds{X}$ hits the attractor in a time of finite mean. Since the attractor is $\mathds{P}$-a.s. of size $\left|A(\omega)\right|=1$, we find
$T_{\Delta}(\omega,x,y)\leq\max\{T_{A}(\omega,x),T_{A}(\omega,y)\}.$
Hence, if the Markov chain is ergodic of degree 2, the synchronization time $T_{\Delta}(\omega,x,y)$ is of finite mean. The next theorem shows that ergodicity of degree 2 is not necessary for this result.
Theorem 5.4.
Let $(\theta,\varphi)$ be a synchronizing Markov RDS representing an irreducible and positive recurrent Markov chain. For any $x,y\in\mathds{X}$, the expected synchronization time is finite, i.e.
$\mathds{E}[T_{\Delta}(\omega,x,y)]<\infty.$
For a proof of this theorem, see Section 5.4. An immediate consequence of this result is that Conjecture 1 holds for synchronizing RDS.
Corollary 5.5.
In a synchronizing RDS, the expected time until attraction $\mathds{E}[T_{A}(\omega,x)]$ is finite for one $x\in\mathds{X}$ if and only if it is finite for all $x\in\mathds{X}$.
Proof.
Assume there is a point $x\in\mathds{X}$ with $\mathds{E}[T_{A}(\omega,x)]<\infty$ and let $y\in\mathds{X}$. We find
$T_{A}(\omega,y)\leq\max\{T_{\Delta}(\omega,x,y),T_{A}(\omega,x)\}.$
By Theorem 5.4, and by assumption, both terms in the maximum have finite expectation. We conclude $\mathds{E}[T_{A}(\omega,y)]<\infty$. ∎
5.3 Examples
We present a class of Markov chains, sometimes referred to as renewal processes, for which one can easily check ergodicity of degree 2 (see e.g. [Mao03]). By Theorem 5.2, this provides examples of RDS for which the expected time to hit the attractor is finite and examples for which the time is infinite.
Example 3.
(Positive recurrent but not ergodic of degree 2) Consider the Markov chain on $\mathds{X}=\mathds{N}_{0}$ with transition probabilities $P(n,n-1)=1$ for $n\in\mathds{N}$ and $P(0,n)=p_{n}$, where $p_{n}>0$ are probabilities that sum to $1$. Note that the Markov chain is irreducible for such $(p_{n})_{n\in\mathds{N}}$. Computing the return time of the state 00,
$\mathds{E}_{0}[\tau_{0}]=\sum_{n\in\mathds{N}}p_{n}(n+1)=1+E_{p},\quad E_{p}:=\sum_{n\in\mathds{N}}np_{n},$
we observe that the Markov chain is positive recurrent if and only if $E_{p}<\infty$.
Furthermore, we compute
$\mathds{E}_{0}[\tau_{0}^{2}]=\sum_{n\in\mathds{N}}p_{n}(n+1)^{2}=1+2E_{p}+E_{p^{2}},\quad E_{p^{2}}:=\sum_{n\in\mathds{N}}n^{2}p_{n}.$
Hence, the Markov chain is ergodic of degree 2 if and only if $E_{p^{2}}<\infty$. In order to construct a Markov chain which is positive recurrent but not ergodic of degree 2, we need to choose the values of $p_{n}$ such that $E_{p}<\infty$ but $E_{p^{2}}=\infty$. An example is $p_{n}=cn^{-\alpha}$, for $\alpha\in(2,3)$, where $c$ is a normalizing constant.
$\Diamond$
The next example provides RDS representations of a Markov chain, which is ergodic of degree 2, with arbitrarily high expected times until attraction.
Example 4.
(Arbitrarily high mean attraction times) Consider the Markov chain on two states $\mathds{X}=\{a,b\}$ with transition probabilities $P(a,a)=P(a,b)=P(b,a)=P(b,b)=\frac{1}{2}$. Given $\epsilon>0$, we construct an RDS representation over a shift space, cf. Section 2.4, such that
$\mathds{P}(\varphi_{\omega}(a,b)$ $=(a,b))=\frac{1-\epsilon}{2},$
$\mathds{P}(\varphi_{\omega}(a,b)$ $=(b,a))=\frac{1-\epsilon}{2},$
$\mathds{P}(\varphi_{\omega}(a,b)$ $=(a,a))=\frac{\epsilon}{2},$
$\mathds{P}(\varphi_{\omega}(a,b)$ $=(b,b))=\frac{\epsilon}{2}.$
At any time, the probability for the two states to coalesce is exactly $\epsilon$. Hence, maximum insulated sets have size $\kappa=1$. The expected time to hit the attractor is geometrically distributed. We have
$\mathds{E}[T_{A}(\omega,a)]=\frac{1}{2}\mathds{E}[T_{A}(\omega,a)\>|\>A(\omega)=b]=\frac{1}{2\epsilon}.$
As $\epsilon\to 0$ this expectation becomes arbitrarily large.
The same technique works for any Markov chain that has an RDS representation with $\kappa>1$. Given an RDS representation with $\kappa>1$, fix $\epsilon>0$ and let all states move independently from one another with probability $\epsilon$. The resulting RDS has $\kappa=1$, but the time until states which were previously insulated coalesce is at least $\frac{1}{\epsilon}$. An interesting open question is whether there is an upper bound on the expected hitting time of the attractor, in particular of the quantity in (28), which is uniform over all RDS representations, in the special case that the Markov chain has no RDS representation with $\kappa>1$. Note that Remark 4.8 provides a sufficient condition for a Markov chain to be not representable by an RDS with $\kappa>1$. $\Diamond$
5.4 Proofs
This section contains the proofs of Theorem 5.2 and Theorem 5.4.
5.4.1 Proof of Theorem 5.2
For the reader’s convenience, we split the proof of the sufficiency of the ergodicity of degree 2 into three steps.
Firstly, assuming that the Markov chain is ergodic of degree 2, we provide an upper bound of $\mathbb{E}[T_{A}(\omega,x)]$ in terms of the expected time for a pair of points $(x,y)$ to either coalesce or become insulated when $x$ is fixed and $y$ is distributed according to $\pi$. This is achieved by constructing a Markov chain on $\mathds{X}^{2}$, which models the two-point motion of the RDS, and considering the hitting time of the set $\Delta\subset\mathds{X}^{2}$ of pairs which have either coalesced, i.e. have reached the diagonal, or are insulated. For technical reasons, which become apparent in the third step, we consider the $N$-step transition probabilities of the two-point Markov chain.
Secondly, we show that the expected hitting time of the set $\Delta$ is finite if the two-point Markov chain satisfies a drift condition outside of a petite set with a Lyapunov function [^1] A formal introduction to $\psi$-irreducibility, petite sets and Lyapunov functions is given in Appendix B. $V:\mathds{X}^{2}\rightarrow[0,\infty)$. Since the one-point Markov chain is ergodic of degree 2, it satisfies a strong drift condition with a potential $W:\mathds{X}\rightarrow[0,\infty)$, which suggests to consider $V(x,y)=W(x)+W(y)$.
Thirdly, we show that the drift condition is indeed fulfilled. This last step is rather technical and is further divided into two substeps.
Afterwards, we proceed to prove that ergodicity of degree 2 is also a necessary condition.
Proof of Theorem 5.2.
A
1. Ergodicity of degree 2 is sufficient.
Step 1.1. Upper bound on $\mathds{E}[T_{A}(\omega,x)]$ in terms of the two-point Markov chain.
Define the set
$\Delta:=\{(x,y)\in\mathds{X}^{2}\>|\>x=y\>\lor\>x\parallel y\}.$
For two points $x,y\in\mathds{X}$, define the random time222 Note that this definition of $T_{\Delta}(\omega,x,y)$ consistently extends the definition (27) for synchronizing RDS.
$T_{\Delta}(\omega,x,y):=\min\{n\in\mathds{N}_{0}\>|\>\varphi^{n}_{\omega}(x,y)\in\Delta\}.$ (29)
Note that $\varphi_{\omega}^{n}(x)\in A(\theta_{n}\omega)$ if and only if all the pairs $\varphi_{\omega}^{n}(x,a)$, for $a\in A(\omega)$, lie in $\Delta$. We find
$T_{A}(\omega,x)=\max_{a\in A(\omega)}T_{\Delta}(\omega,x,a).$
This enables us to estimate the expectation of $T_{A}(\omega,x)$ via considering the times $T_{\Delta}(\omega,x,y)$ for pairs $(x,y)\in\mathds{X}^{2}$:
$\mathds{E}[T_{A}(\omega,x)]$ $\leq\mathds{E}\left[\sum_{a\in A(\omega)}T_{\Delta}(\omega,x,a)\right]=\mathds{E}\left[\sum_{y\in\mathds{X}}\mathds{1}_{A(\omega)}(y)T_{\Delta}(\omega,x,y)\right]$
$=\sum_{y\in\mathds{X}}\mathds{E}[\mathds{1}_{A(\omega)}(y)T_{\Delta}(\omega,x,y)].$
Note that all quantities are non-negative such that the exchange of an infinite sum and the expectation is justified. Since $\mathds{1}_{A(\omega)}$ is $\mathcal{F}_{-\infty}^{0}$-measurable and $T_{\Delta}(\omega,x,y)$ is $\mathcal{F}_{0}^{\infty}$-measurable, we obtain
$\mathds{E}[T_{A}(\omega,x)]$ $\leq\sum_{y\in\mathds{X}}\mathds{E}[\mathds{1}_{A(\omega)}(y)]\mathds{E}[T_{\Delta}(\omega,x,y)]$ (30)
$=\sum_{y\in\mathds{X}}\kappa\pi(y)\mathds{E}[T_{\Delta}(\omega,x,y)],$
recalling that $\mathds{P}(y\in A(\omega))=\kappa\pi(y)$, cf. (13). We now define a Markov chain on the state space $\mathcal{S}:=\mathds{X}^{2}\setminus\Delta\cup\{\partial\}$, in which all pairs $(x,y)\in\Delta$ are identified as a single state denoted by $\partial$, via the transition probabilities
$Q((x,y),(x^{\prime},y^{\prime}))$ $=\mathds{P}(\varphi_{\omega}(x,y)=(x^{\prime},y^{\prime})\},$ $(x,y),(x^{\prime},y^{\prime})$ $\in\mathds{X}^{2}\setminus\partial,$ (31)
$Q((x,y),\partial)$ $=\mathds{P}(\varphi_{\omega}(x,y)\in\Delta\},$ $(x,y)$ $\in\mathds{X}^{2}\setminus\partial,$
$Q(\partial,\partial)$ $=1.$
Note that $\partial$ is the unique, absorbing state of the Markov chain. In particular, the value $\mathds{E}[T_{\Delta}(\omega,x,y)]$ equals the expected hitting time of the state $\partial\in\mathcal{S}$ for the initial state $(x,y)\in\mathcal{S}^{2}\setminus\partial$ under the new Markov chain defined via (31), i.e.
$\mathds{E}[T_{\Delta}(\omega,x,y)]=\mathds{E}^{Q}_{(x,y)}[\tau_{\partial}],$ (32)
where $\mathds{E}^{Q}$ refers to the expectation of the Markov chain with transition probabilities $Q$. Choose some $N\in\mathbb{N}$ and define the $N$-step transition probabilities
$R(s,s^{\prime})=Q^{N}(s,s^{\prime}).$
where $s,s^{\prime}\in\mathcal{S}$. This defines transition probabilities for another Markov chain on $\mathcal{S}$ for which $\partial$ is the unique, absorbing set as well. We clearly find
$\mathds{E}_{(x,y)}^{Q}[\tau_{\partial}]\leq N\mathds{E}_{(x,y)}^{R}[\tau_{\partial}].$ (33)
Inserting this inequality and (32) into (30) yields
$\mathds{E}[T_{A}(\omega,x)]\leq N\kappa\sum_{y\in\mathds{X}}\pi(y)\mathds{E}_{(x,y)}^{R}[\tau_{\partial}]$ (34)
This bound concludes step 1.1. $\Diamond$
Fix a base point $x_{0}\in\mathds{X}$. For $x\in\mathds{X}$, set $W(x):=\mathds{E}_{x}[\tau_{x_{0}}]$ with the exception $W(x_{0}):=0$. Define a Lyapunov function $V:\mathcal{S}\to[0,\infty)$ by
$V(x,y):=W(x)+W(y),\qquad V(\partial):=0.$ (35)
In the next step, we show that the right hand side of (34) is finite if the drift condition
$RV(s)\leq V(s)-1+b\mathds{1}_{C}(s),\qquad\forall s\in\mathcal{S}$ (36)
is satisfied, where $b>0$ is a constant, $C\subset\mathcal{S}$ is a petite set with respect to $R$, and $RV$ is defined by [^3] In particular, we have $RV(\partial)=0$
$RV(s)=\sum_{s^{\prime}\in\mathcal{S}}R(s,s^{\prime})V(s^{\prime}),\ s\in\mathcal{S}.$ (37)
A brief, but formal introduction to Lyapunov functions and petite sets is included in Appendix B. For our purposes, the following two statements about petite sets are sufficient:
  1. (i)
    Let $C\subset\mathcal{S}$ such that there is a state $s^{\prime}\in\mathcal{S}$ and $\delta>0$ such that for all $s\in C$ we find $R(s,s^{\prime})>\delta$. Then, $C$ is petite with respect to $R$;
  2. (ii)
    The union of two petite sets is petite.
Step 1.2. $\mathbb{E}[T_{A}(\omega,x)]<\infty$ for all $x\in\mathds{X}$ if the drift condition (36) holds.
By [MT12, Theorem 11.3.11], see also Theorem B.4, condition (36) implies that there is a constant $c(\partial)$ such that for each $s\in\mathcal{S}\setminus\partial$,
$\mathds{E}_{s}^{R}[\tau_{\partial}]\leq V(s)+c(\partial).$ (38)
Assume for now that $V$, as defined in (35), satisfies the drift condition (36) for a certain choice of $N\in\mathds{N}$, $b>0$, and petite set $C$. Using (34), this yields
$\mathds{E}[T_{A}(\omega,x)]$ $\leq N\kappa\sum_{y\in\mathds{X}}\pi(y)\mathds{E}_{(x,y)}^{R}[\tau_{\partial}]$
$\leq N\kappa\sum_{y\in\mathds{X}}\pi(y)(V(x,y)+c(\partial))$
$=N\kappa\left(c(\partial)+W(x)+\sum_{y\in\mathds{X}}\pi(y)W(y)\right).$
By definition of $W$, we have
$\sum_{y\in\mathds{X}}\pi(y)W(y)=\mathds{E}_{\pi}[\tau_{x_{0}}]-\mathds{E}_{x_{0}}[\tau_{x_{0}}].$
This quantity is finite by the assumption of ergodicity of degree 2; see (iii) in Proposition 5.1. We conclude that $\mathds{E}[T_{A}(\omega,x)]$ is finite for all $x\in\mathds{X}$ if the drift condition (36) is satisfied.
As a direct consequence, we obtain the stronger result which reads
$\sum_{x\in\mathds{X}}\pi(x)\mathds{E}[T_{A}(\omega,x)]$ $\leq N\kappa\left(c(\partial)+\sum_{x\in\mathds{X}}\pi(x)W(x)+\sum_{y\in\mathds{X}}\pi(y)W(y)\right)$
$=N\kappa\left(c(\partial)+2\mathds{E}_{\pi}[\tau_{x_{0}}]-2\mathds{E}_{x_{0}}[\tau_{x_{0}}]\right)$
<<$\infty$.< ∞ .
This completes Step 1.2. $\Diamond$
It remains to show that the drift condition of $V$ is indeed satisfied for some $N\in\mathbb{N}$, $b>0$, and a petite set $C\subset\mathds{X}$.
Step 1.3. Verification of the drift condition (36).
For $(x,y)\in\mathcal{S}\setminus\partial$, we have
$RV(x,y)$ $=\sum_{(x^{\prime},y^{\prime})\in\mathcal{S}\setminus\partial}R((x,y),(x^{\prime},y^{\prime}))(W(x^{\prime})+W(y^{\prime}))$ (39)
$=\sum_{(x^{\prime},y^{\prime})\in\mathcal{S}\setminus\partial}Q^{N}((x,y),(x^{\prime},y^{\prime}))(W(x^{\prime})+W(y^{\prime})).$
Since $\partial$ is an $\mathds{P}$-a.s. absorbing set of the RDS, we have for $(x,y),(x^{\prime},y^{\prime})\in\mathcal{S}\setminus\partial$
$Q^{N}((x,y),(x^{\prime},y^{\prime}))=\mathds{P}(\varphi^{N}_{\omega}(x,y)=(x^{\prime},y^{\prime})).$
We compute
$\sum_{(x^{\prime},y^{\prime})\in\mathcal{S}\setminus\partial}Q^{N}((x,y),(x^{\prime},y^{\prime}))W(x^{\prime})$ $=\sum_{x^{\prime}\in\mathds{X}}\sum_{\begin{subarray}{c}y^{\prime}\in\mathds{X},\>\textup{s.t.}\\ (x^{\prime},y^{\prime})\in\mathcal{S}\setminus\partial\end{subarray}}\mathds{P}(\varphi^{N}_{\omega}(x,y)=(x^{\prime},y^{\prime}))W(x^{\prime})$
$\leq\sum_{x^{\prime}\in\mathds{X}}\sum_{y^{\prime}\in\mathds{X}}\mathds{P}(\varphi^{N}_{\omega}(x,y)=(x^{\prime},y^{\prime}))W(x^{\prime})$
$=\sum_{x^{\prime}\in\mathds{X}}\mathds{P}(\varphi^{N}_{\omega}(x)=x^{\prime})W(x^{\prime})$
$=P^{N}W(x),$
where $P^{N}W$ is defined analogously to (37). Inserting this estimate and the analogous one for $P^{N}W(y)$ into (39) yields
$RV(x,y)\leq\left(P^{N}W(x)+P^{N}W(y)\right).$ (40)
To determine $P^{N}W(x)$, we compute the values $P^{n}W(x)$ recursively. For $n=1$, we have
$PW(x)=\sum_{x^{\prime}\in\mathds{X}}P(x,x^{\prime})W(x^{\prime})=\sum_{x^{\prime}\neq 0}P(x,x^{\prime})\mathds{E}_{x^{\prime}}[\tau_{x_{0}}]=\mathds{E}_{x}[\tau_{x_{0}}]-1,$
which can be written as
$PW(x)=\begin{cases}W(x)-1,\quad&x\neq x_{0},\\ \mathds{E}_{x_{0}}[\tau_{x_{0}}]-1,&x=x_{0}.\end{cases}$
We will simply write $PW(x_{0})$ instead of $\mathds{E}_{x_{0}}[\tau_{x_{0}}]-1$. Proceeding recursively in $n$, we obtain
$P^{n+1}W(x)$ $=\sum_{x^{\prime\prime}\in\mathds{X}}P^{n+1}(x,x^{\prime\prime})W(x^{\prime\prime})=\sum_{x^{\prime\prime}\in\mathds{X}}\sum_{x^{\prime}\in\mathds{X}}P^{n}(x,x^{\prime})P(x^{\prime},x^{\prime\prime})W(x^{\prime\prime})$
$=\sum_{x^{\prime}\in\mathds{X}}P^{n}(x,x^{\prime})PW(x^{\prime})=P^{n}(x,x_{0})PW(x_{0})+\sum_{x^{\prime}\neq x_{0}}P^{n}(x,x^{\prime})(W(x^{\prime})-1)$
$=P^{n}(x,x_{0})PW(x_{0})+P^{n}W(x)-1+P^{n}(x,x_{0})$
$=P^{n}W(x)-1+P^{n}(x,x_{0})\mathds{E}_{x_{0}}[\tau_{x_{0}}].$
For $x\neq x_{0}$, this recursion has the simple solution
$P^{n}W(x)=W(x)-n+\sum_{m=1}^{n-1}P^{m}(x,x_{0})\mathds{E}_{x_{0}}[\tau_{x_{0}}].$
Recalling the identity $\mathds{E}_{x_{0}}[\tau_{x_{0}}]=\frac{1}{\pi(x_{0})}$ from Theorem 2.1, we can write equivalently
$P^{n}W(x)=W(x)-1+\mathds{E}_{x_{0}}[\tau_{x_{0}}]\sum_{m=1}^{n-1}\big{(}P^{m}(x,x_{0})-\pi(x_{0})\big{)}.$
Inserting this formula into (40) yields
$RV(x,y)$ $\leq W(x)+W(y)-2+\mathds{E}_{x_{0}}[\tau_{x_{0}}]\sum_{n=1}^{N-1}\big{(}P^{n}(x,x_{0})-\pi(x_{0})+P^{n}(y,x_{0})-\pi(x_{0})\big{)}$ (41)
$\leq V(x,y)-1+\mathds{E}_{x_{0}}[\tau_{x_{0}}]\sum_{n=1}^{N-1}\big{(}P^{n}(x,x_{0})-\pi(x_{0})+P^{n}(y,x_{0})-\pi(x_{0})\big{)}.$
In the following, we will determine for which pairs of points $(x,y)\in\mathcal{S}\setminus\partial$, the last term is non-positive.
Step 1.3.1: There is a constant $K>0$ such that for any $x\in\mathds{X},\,N\in\mathbb{N}$,
$\sum_{n=1}^{N-1}\big{(}P^{n}(x,x_{0})-\pi(x_{0})\big{)}\leq K.$ (42)
Let $\psi_{n}(x):=\mathds{P}_{x}(\tau(x_{0})=n)$ be the probability that the Markov chain $(X_{n})_{n\in\mathds{N}}$, initialized at $X_{0}=x$, visits the state $x_{0}$ at time $n$ for the first time. For $x=x_{0}$, this means the first return to $x_{0}$ after at least one step. We have the two identities
$\sum_{n=1}^{\infty}\psi_{n}(x)=1,\qquad P^{n}(x,x_{0})=\sum_{m=1}^{n}\psi_{m}(x)P^{n-m}(x_{0},x_{0}),$
for any $x\in\mathds{X}$, where we use the convention $P^{0}(x_{0},x_{0})=1$. We compute
$\sum_{n=1}^{N-1}\big{(}P^{n}(x,x_{0})-\pi(x_{0})\big{)}$ $=\sum_{n=1}^{N-1}\sum_{m=1}^{n}\psi_{m}(x)P^{n-m}(x_{0},x_{0})-(N-1)\pi(x_{0})$ (43)
$\leq\sum_{m=1}^{N-1}\psi_{m}(x)\sum_{n=0}^{N-1-m}P^{n}(x_{0},x_{0})-\sum_{m=1}^{N-1}\psi_{m}(x)(N-1)\pi(x_{0})$
$\leq\sum_{m=1}^{N-1}\psi_{m}(x)\sum_{n=0}^{N-1-m}\big{(}P^{n}(x_{0},x_{0})-\pi(x_{0})\big{)}.$
As we assumed the Markov chain on $\mathds{X}$ with transition probabilities $P$ to be ergodic of degree 2, we can deduce from [Mao03, Theorem 1.1] that the deviation matrix
$d_{x,y}:=\sum_{n=1}^{\infty}\big{(}P^{n}(x,y)-\pi(y)\big{)}$
is finite for any $x,y\in\mathds{X}$. In particular, $d_{x_{0},x_{0}}$ is a well-defined finite constant. Hence, there is a constant $K>0$ such that
$\sum_{m=1}^{n}\big{(}P^{m}(x_{0},x_{0})-\pi(x_{0})\big{)}\leq K$
for any $n\in\mathds{N}$, which, inserted into estimate (43), yields the claim (42).
Step 1.3.2: Construction of a petite set $C$.
We introduce
$\overline{C}_{N}$ :=$:=\left\{x\in\mathds{X}\>\Bigg{|}\>\sum_{n=1}^{N-1}\big{(}P^{n}(x,x_{0})-\pi(x_{0})\big{)}\geq-K\right\},$
$C_{N}$ :=$:=\big{\{}(x,y)\in\mathcal{S}\setminus\partial\>|\>x\in\overline{C}_{N}\land y\in\overline{C}_{N}\big{\}}\cup\{\partial\}.$
Define the constant $b:=2\mathds{E}_{0}[\tau_{0}]K$. By choosing $K\geq\frac{1}{2}$, we assume w.l.o.g. $b\geq 1$. We show that, for all $s\in\mathcal{S}$ and $N\in\mathbb{N}$, $V$ satisfies the drift condition
$RV(s)\leq V(s)-1+b\mathds{1}_{C_{N}}(s).$ (44)
  • $s=\partial$: the inequality holds since $RV(\partial)=V(\partial)=0$ and $\partial\in C_{N}$.
  • $(x,y)\in\mathcal{S}\setminus C_{N}$: without loss of generality, let $x\notin\overline{C}_{N}$. Using (42), equation (41) yields
    $RV(x,y)$ $\leq V(x,y)-1+\mathds{E}_{x_{0}}[\tau_{x_{0}}]\left(\sum_{n=1}^{N-1}\big{(}P^{n}(x,x_{0})-\pi(x_{0})\big{)}+K\right)$
    $\leq V(x,y)-1\leq V(x,y)-1+b\mathds{1}_{C_{N}}(x,y).$
  • $(x,y)\in C_{N}\setminus\partial$: again, using (42), we obtain
    $RV(x,y)\leq V(x,y)-1+2\mathds{E}_{x_{0}}[\tau_{x_{0}}]K=V(x,y)-1+b\mathds{1}_{C_{N}}(x,y).$
It remains to choose $N\in\mathds{N}$ appropriately, such that $C_{N}$ is petite with respect to $R$. Since any singleton set is petite, we only need to show that $C_{N}\setminus\partial$ is petite. For $x\in\overline{C}_{N}$, we compute
$(N-1)\pi(x_{0})-K$ $\leq\sum_{n=1}^{N-1}P^{n}(x,x_{0})=\sum_{m=1}^{N-1}\psi_{m}(x)\sum_{n=0}^{N-1-m}P^{n}(x_{0},x_{0})$
$=\sum_{m=1}^{N-1}\psi_{m}(x)\left(1+(N-1-m)\pi(x_{0})+\sum_{n=1}^{N-1-m}\big{(}P^{n}(x_{0},x_{0})-\pi(x_{0})\big{)}\right)$
$\leq\sum_{m=1}^{N-1}\psi_{m}(x)\left(1+(N-1)\pi(x_{0})+K\right),$
which gives
$\sum_{m=1}^{N-1}\psi_{m}(x)\geq\frac{(N-1)\pi(x_{0})-K}{(N-1)\pi(x_{0})+K+1}.$ (45)
Note that for $N\to\infty$ the right-hand side converges to $1$. Hence, for $0<\epsilon<\frac{1}{2}$, we can choose $N\in\mathds{N}$ such that
$\frac{(N-1)\pi(x_{0})-K}{(N-1)\pi(x_{0})+K+1}>1-\epsilon>\frac{1}{2}.$
We can now set $C:=C_{N}$ for this choice of $N\in\mathbb{N}$. Let $(x,y)\in C\setminus\Delta$. By (45), the probability that $x$, respective $y$, visits the state $x_{0}$ within $N-1$ steps is greater than $1-\epsilon$. We find
$\mathds{P}(\exists k,\ell<N:\>\varphi_{\omega}^{k}(x)=x_{0}\land\varphi_{\omega}^{\ell}(y)=x_{0})>1-2\epsilon>0.$
In particular, for any $(x,y)\in C\setminus\Delta$ there are numbers $k,\ell\leq N-1$ such that
$\mathds{P}\big{(}\varphi_{\omega}^{k}(x)=x_{0}\land\varphi_{\omega}^{\ell}(x)=x_{0}\big{)}>\frac{1-2\epsilon}{(N-1)^{2}}=:\delta>0.$
For $k,\ell\leq N-1$, we define
$C^{(k,\ell)}:=\{(x,y)\in C\setminus\Delta\>|\>\mathds{P}(\varphi_{\omega}^{k}(x)=x_{0}\land\varphi_{\omega}^{\ell}(x)=x_{0})>\delta\}\subset C\setminus\{\Delta\},$
and choose finite sets $\chi_{k},\chi_{\ell}\subset\mathds{X}$ such that $P^{N-k}(0,\chi_{k})>1-\frac{\delta}{3}$ and $P^{N-\ell}(0,\chi_{\ell})>1-\frac{\delta}{3}$. Hence, for any $(x,y)\in C^{(k,\ell)}$ we find
$\mathds{P}\big{(}\varphi_{\omega}^{N}(x)\in\chi_{k}\land\varphi_{\omega}^{N}(x)\in\chi_{\ell}\big{)}>\frac{\delta}{3}.$
Let $M_{k,\ell}:=\left|\chi_{k}\times\chi_{\ell}\right|<\infty$. Then for each $(x,y)\in C^{(k,\ell)}$ there is a pair $(x^{\prime},y^{\prime})\in\chi_{k}\times\chi_{\ell}$ such that $P(\varphi_{\omega}^{N}(x,y)=(x^{\prime},y^{\prime}))>\frac{\delta}{3M_{k,\ell}}$ and, hence, the union of the sets
$C^{(k,\ell)}_{(x^{\prime},y^{\prime})}:=\left\{(x,y)\in C\setminus\Delta\>\Big{|}\>\mathds{P}(\varphi_{\omega}^{N}(x,y)=(x^{\prime},y^{\prime})>\frac{\delta}{3M_{k,\ell}}\right\},\ k,\ell\leq N-1,(x^{\prime},y^{\prime})\in\chi_{k}\times\chi_{\ell},$
equals $C\setminus\{\Delta\}$. Interpreting each pair $(x^{\prime},y^{\prime})$ as a state $s^{\prime}\in\mathcal{S}$, we obtain
$R(s,s^{\prime})>\frac{\delta}{3M_{k,\ell}},\qquad\forall s\in C^{(k,\ell)}_{(x^{\prime},y^{\prime})},$
which shows that the sets $C^{(k,\ell)}_{(x^{\prime},y^{\prime})}$ are petite with respect to $R$. Since the finite union of petite sets is petite, we can conclude that $C$ is petite. This was the last remaining step of the proof that ergodicity of degree 2 is sufficient. $\Diamond$
2. Ergodicity of degree 2 is necessary.
Recall from Section 4.3 the Markov chain on the set of maximum insulated sets $\mathcal{I}$, with transition probabilities
$S(K,L)=\mathds{P}(\varphi_{\omega}(K)=L),$
for $K,L\in\mathcal{I}$, and the set $\mathcal{A}$ of maximum insulated sets which the attractor admits with positive probability, cf. (23). According to Theorem 4.9, $\mathcal{A}$ is the unique irreducibility class of the Markov chain on $\mathcal{I}$, with stationary distribution
$\varrho(K)=\mathds{P}(A(\omega)=K),\quad K\in\mathcal{A}.$
Let $\overline{\Omega}\subset\Omega$ be the $\theta$-invariant subset of full probability such that (24) holds. Restricted to $\overline{\Omega}$, the RDS $(\theta,\varphi)$ is an RDS representation of the Markov chain defined by $S$.
Assume now that the expected time $\mathds{E}[T_{A}(\omega,x)]$ is finite for all $x\in\mathds{X}$. We fix some maximum insulated set $K\in\mathcal{A}$ and define the hitting time
$T_{A}(\omega,K)=\min\{n\in\mathds{N}\>|\>\varphi_{\omega}^{n}(K)=A(\theta_{n}\omega)\}.$
We can deduce from the assumption that
$\mathds{E}[T_{A}(\omega,K)]=\mathds{E}\left[\max_{x\in K}T_{A}(\omega,x)\right]<\infty.$
Let $L\in\mathcal{A}$ and observe that
$\mathds{P}(T_{A}(\omega,K)>n)\geq\left|\mathds{P}(\varphi^{n}_{\omega}(K)=L)-\mathds{P}(A(\theta_{n}\omega)=L)\right|$ $=\left|S^{n}(K,L)-\varrho(L)\right|,$
which yields
$\sum_{n\in\mathds{N}}\left|S^{n}(K,L)-\varrho(L)\right|$ $\leq\sum_{n\in\mathds{N}}\mathds{P}(T_{A}(\omega,K)>n)=\mathds{E}[T_{A}(\omega,K)]-1<\infty.$
Hence, the sum converges absolutely and the deviation matrix $d_{K,L}$ is well-defined and finite
$d_{K,L}:=\sum_{n\in\mathds{N}}\big{(}S^{n}(K,L)-\varrho(L)\big{)}<\infty.$
By [Mao03, Theorem 1.1], this means that the Markov chain with transition probabilities $S$, restricted to $\mathcal{A}$, is ergodic of degree 2.
Fix $K_{0}\in\mathcal{A}$ and, for $K\in\mathcal{A}$, define the hitting time
$\tau_{K_{0}}(\omega,K):=\min\{n\in\mathds{N}\>|\>\varphi_{\omega}^{n}(K)=K_{0}\}.$
By Proposition 5.1, we have $\mathds{E}_{\varrho}^{S}[\tau_{K_{0}}]<\infty$, and we compute
$\infty>\mathds{E}_{\varrho}^{S}[\tau_{K_{0}}]$ $=\sum_{K\in\mathcal{A}}\varrho(K)\mathds{E}_{K}^{S}[\tau_{K_{0}}]=\sum_{K\in\mathcal{A}}\varrho(K)\mathds{E}[\tau_{K_{0}}(\omega,K)]$ (46)
$=\sum_{K\in\mathcal{A}}\mathds{P}(A(\omega)=K)\>\mathds{E}[\tau_{K_{0}}(\omega,A(\omega))\>|\>A(\omega)=K]$
$=\mathds{E}[\tau_{K_{0}}(\omega,A(\omega))].$
Recall that $\mathds{P}$-a.s. the set $A(\omega)$ is of size $\kappa$ and observe that
[…]
$=\frac{1}{\kappa}\sum_{x\in\mathds{X}}\mathds{P}(x\in A(\omega))\>\mathds{E}[\tau_{K_{0}}(\omega,A(\omega))\>|\>x\in A(\omega)],$
$=\sum_{x\in\mathds{X}}\pi(x)\mathds{E}[\tau_{K_{0}}(\omega,A(\omega))\>|\>x\in A(\omega)],$
where we used again that the probability for a point $x\in\mathds{X}$ to lie in $A(\omega)$ is $\kappa\pi(x)$, cf. (13). Given $x\in A(\omega)$, the time for $A(\omega)$ to reach the set $K_{0}$ is at least as long as the time for $x$ to reach $K_{0}$. By abuse of notation, we denote by $\tau_{K_{0}}$ the random time it takes for a single point $x\in\mathds{X}$ to hit the set $K_{0}$. We get the bound
$\sum_{x\in\mathds{X}}\pi(x)\mathds{E}[\tau_{K_{0}}(\omega,A(\omega))\>|\>x\in A(\omega)]$ $\geq\sum_{x\in\mathds{X}}\pi(x)\mathds{E}[\tau_{K_{0}}(\omega,x)]$ (48)
$=\sum_{x\in\mathds{X}}\pi(x)\mathds{E}_{x}[\tau_{K_{0}}]$
$=\mathds{E}_{\pi}[\tau_{K_{0}}].$
Combining (46)--(48) yields
$\mathds{E}_{\pi}[\tau_{K_{0}}]\leq\mathds{E}_{\varrho}^{S}[\tau_{K_{0}}]<\infty.$
For any $y\in\mathds{X}$, we can now conclude that
$\mathds{E}_{\pi}[\tau_{y}]\leq\mathds{E}_{\pi}[\tau_{K_{0}}]+\max_{x\in K_{0}}\mathds{E}_{x}[\tau_{y}]<\infty,$
which shows the claim by Proposition 5.1. ∎
5.4.2 Proof of Theorem 5.2
Let $P$ be the transition probabilities of the associated Markov chain and denote its stationary distribution by $\pi$.
Consider the Markov chain on $\mathds{X}^{2}$ which models the 2-point motion of the RDS $(\theta,\varphi)$. Define the transition probabilities
$Q((x,y),(x^{\prime},y^{\prime})):=\mathds{P}(\varphi_{\omega}(x,y)=(x^{\prime},y^{\prime})).$
Note that the diagonal
$\Delta:=\{(x,x)\in\mathds{X}^{2}\mid x\in\mathds{X}\},$
is the unique irreducibility class of this Markov chain. To distinguish probabilities and expected values of the Markov chain on $\mathds{X}^{2}$ defined by $Q$ from those of the Markov chain on $\mathds{X}$ defined by $P$, we write $\mathds{P}^{Q}$ and $\mathds{E}^{Q}$. Since $Q$ models the 2-point motion of the RDS, we find
$\mathds{E}[T_{\Delta}(\omega,x^{*},y^{*})]=\mathds{E}_{(x^{*},y^{*})}^{Q}[\tau_{\Delta}],$ (49)
for any $x^{*},y^{*}\in\mathds{X}$, where $\tau_{\Delta}$ is the hitting time of the set $\Delta$.
We provide a sketch of the proof idea. We define a set $\Gamma\subset\mathds{X}^{2}$ which connects $\Delta$ to $(x^{*},y^{*})$ and show that the Markov chain defined by $Q$ hits $\Gamma\cup\Delta$ in a time of finite mean, when started in $(x^{*},y^{*})$. Hence, after time $\tau_{\Delta\cup\Gamma}$, the Markov chain has either hit the diagonal $\Delta$, or has landed in a point $(x^{**},y^{**})$, which is ’closer’ to $\Delta$ than $(x^{*},y^{*})$. We then proceed inductively and show that in each step the set $\Gamma$ decreases until $\Gamma=\emptyset$, and we are guaranteed to have hit the diagonal $\Delta$. To show that the Markov chain defined by $Q$ hits the set $\Delta\cup\Gamma$ in a time of finite mean, we alter the transition probabilities $Q$ inside of the set $\Gamma$, resulting in transition probabilities $R$, such that $(x^{*},y^{*})$ lies in the unique irreducibility class $I$ of the Markov chain defined by $R$. We then show that the Markov chain defined by $R$ is positive recurrent on $I$ which shows that the set $\Delta\cup\Gamma$ is hit in a time of finite mean. See Figure 2 for a sketch of the setting.
Proof of Theorem 5.2.
Let $x^{*},y^{*}\in\mathds{X}$. A path in $\mathds{X}$ is a sequence $\gamma=(a_{1},\ldots,a_{k})$ of elements $a_{i}\in\mathds{X}$ such that $P(a_{i},a_{i+1})>0$ for all $11\leq i\leq k-$ 1. Since the associated Markov chain is irreducible, for each $z\in\mathds{X}$, there are paths $\gamma_{x^{*}}$ and $\gamma_{y^{*}}$ in $\mathds{X}$ that start in $z$ and end in $x^{*}$ and $y^{*}$, respectively. We write these paths as $\gamma_{x^{*}}=(a_{1},\ldots,a_{k})$ and $\gamma_{y^{*}}=(b_{1},\ldots,b_{\ell})$ with $a_{1}=b_{1}=z$, $a_{k}=x^{*}$ and $b_{\ell}=y^{*}$. Since the RDS $(\theta,\varphi)$ is synchronizing, the associated Markov chain is aperiodic, cf. Remark 4.6. Hence, for each $z\in\mathds{X}$, there are paths $\gamma_{x^{*}}$ and $\gamma_{y^{*}}$ of the same length, i.e. $k=\ell$, cf. [BH22, Section 2.4]. Choose $z\in\mathds{X}$ and paths $\gamma_{x^{*}}$, $\gamma_{y^{*}}$ such that $k=\ell$ is minimal. Define the set
$\Gamma:=\{(a_{i},b_{i})\in\mathds{X}^{2}\mid 1\leq i\leq k-1\}\subset\mathds{X}^{2}.$
Since the paths $\gamma_{x^{*}}$ and $\gamma_{y^{*}}$ are chosen to be minimal, the intersection of $\Delta$ and $\Gamma$ is $\{(z,z)\}$.
We now show that $\mathds{E}_{(x^{*},y^{*})}^{Q}[\tau_{\Delta\cup\Gamma}]<\infty$, where $\tau_{\Delta\cup\Gamma}$ is the hitting time of the Markov chain defined by $Q$ of the set $\Delta\cup\Gamma$. Define another Markov chain on $\mathds{X}^{2}$ that is equivalent to the one defined by $Q$ outside of the set $\Gamma$. Inside of $\Gamma$, the two components move independently from one another. Define
$R((x,y),(x^{\prime},y^{\prime})):=\begin{cases}Q((x,y),(x^{\prime},y^{\prime})),\quad&(x,y)\notin\Gamma,\\ P(x,x^{\prime})P(y,y^{\prime}),\quad&(x,y)\in\Gamma.\end{cases}$ (50)
Since, outside of $\Gamma$, the transition probabilities $R$ and $Q$ coincide, we find
$\mathds{E}_{(x^{*},y^{*})}^{Q}[\tau_{\Delta\cup\Gamma}]=\mathds{E}_{(x^{*},y^{*})}^{R}[\tau_{\Delta\cup\Gamma}].$ (51)
Unlike for the Markov chain defined by $Q$, the set $\Delta$ is not an irreducibility class of the Markov chain defined by $R$, since the diagonal can be exited in the point $(z,z)\in\Gamma$. In particular, the point $(x^{*},y^{*})$ can be reached from the diagonal with positive probability. Let $I$ be the unique irreducibility class of the Markov chain defined by $R$. We have $\Delta\cup\Gamma\subset I$ and, in particular, $(x^{*},y^{*})\in I$. We show that the Markov chain with transition probabilities $R$ is positive recurrent when restricted to $I$.
We denote the Markov chain on $\mathds{X}^{2}$ with transition probabilities $R$ by $(\mathbf{X}_{n})_{n\in\mathds{N}}$. We consider the two components $\mathbf{X}_{n}^{1}$ and $\mathbf{X}_{n}^{2}$. Note that each of these two components is a Markov chain on $\mathds{X}$ with transition probabilities $P$, since
$\sum_{y^{\prime}\in\mathds{X}}R((x,y),(x^{\prime},y^{\prime}))=P(x,x^{\prime}),$
for any $x,x^{\prime}\in\mathds{X}$ and $y\in\mathds{X}$. Both of the the Markov chains $(\mathbf{X}^{1}_{n})_{n\in\mathds{N}}$ and $(\mathbf{X}^{2}_{n})_{n\in\mathds{N}}$ are known to be positive recurrent on $\mathds{X}$ with stationary distribution $\pi$. Let $C\subset\mathds{X}$ be a finite set with $\pi(C)>\frac{1}{2}$. Using the Birkhoff ergodic theorem for Markov chains, cf. Theorem 2.1, we compute
$\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}\mathds{1}_{C^{2}}(\mathbf{X}_{n})=$ $\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}1-\mathds{1}_{\mathds{X}^{2}\setminus C^{2}}(\mathbf{X}_{n})$
$\geq$ $\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}1-\mathds{1}_{\mathds{X}\setminus C}(\mathbf{X}_{n}^{1})-\mathds{1}_{\mathds{X}\setminus C}(\mathbf{X}_{n}^{2})$
$=$ $1-\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}\mathds{1}_{\mathds{X}\setminus C}(\mathbf{X}_{n}^{1})-\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}\mathds{1}_{\mathds{X}\setminus C}(\mathbf{X}_{n}^{2})$
$=$ $0\displaystyle 1-2(1-\pi(C))>$ 0 .
The set $C^{2}$ is finite. Hence, there is at least one pair $\mathbf{c}\in C^{2}$ such that
$\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}\mathds{1}_{\mathbf{c}}(\mathbf{X}_{n})>0.$ (52)
If we choose $\mathbf{X}_{0}\in I$, then $\mathbf{X}_{n}\in I$ for all $n\in\mathds{N}$ since $I$ is an irreducibility class. Hence, there is a $\mathbf{c}\in I$ such that the limit in (52) is positive. This shows that the Markov chain $(\mathbf{X}_{n})_{n\in\mathds{N}}$, restricted to $I$, is recurrent. A classic result from Markov chain theory, e.g. [BH22, Proposition 2.1], states that for an irreducible, recurrent Markov chain $(\mathbf{X}_{n})_{n\in\mathds{N}}$ on $I$ and $\mathbf{c}\in I$, we find the following ($\mathds{P}^{R}_{(x,y)}$-a.s. limit for any initial condition $(x,y)\in I$,
$\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}\mathds{1}_{\mathbf{c}}(\mathbf{X}_{n})=\frac{1}{\mathds{E}_{\mathbf{c}}[\tau_{\mathbf{c}}]}.$ (53)
We emphasize that this result only requires recurrence of $(\mathbf{X}_{n})_{n\in\mathds{N}}$ and not positive recurrence. In particular, the Markov chain is positive recurrent if and only if the quantities in (53) are positive. In (52) we showed that the limit on the left-hand side is positive. This shows that the Markov chain $(\mathbf{X}_{n})_{n\in\mathds{N}}$ with transition probabilities $R$ is positive recurrent when restricted to $I$. This implies that $\mathds{E}^{R}_{(x,y)}[\tau_{\Delta\cup\Gamma}]<\infty$ for all $(x,y)\in I$. For the starting point $(x^{*},y^{*})$, equation (51) yields $\mathds{E}_{(x^{*},y^{*})}^{Q}[\tau_{\Delta\cup\Gamma}]<\infty$.
After time $\tau_{\Delta\cup\Gamma}$, the Markov chain defined by $Q$ and started in $(x^{*},y^{*})$ has either hit $\Delta$ or has entered the set $\Gamma$ in a point $(a_{i},b_{i})$ for $i\geq 2$. In the latter case, there are paths $\gamma_{a_{i}}=(a_{1},\ldots,a_{i})$ and $\gamma_{b_{i}}=(b_{1},\ldots,b_{i})$ from $a_{1}=b_{1}=z$ to $a_{i}$ and $b_{i}$ respectively. Both paths are of length $i$, which is strictly less than $k$. We proceed inductively, with the new initial point $(a_{i},b_{i})$. After at most $k$ iterations, we find $(a_{i},b_{i})\in\Delta$. Since each step is completed in a time of finite mean, we conclude $\mathds{E}_{(x^{*},y^{*})}^{Q}[\tau_{\Delta}]<\infty$ and by (49) we have $\mathds{E}[T_{\Delta}(\omega,x^{*},y^{*})]<\infty$. This completes the proof. ∎
[…]
Figure 2: A schematic representation of the state space $\mathds{X}^{2}$. The diagonal set $\Delta$ is marked as a dashed line. The set $\Gamma$ connects $\Delta$ to the point $(x^{*},y^{*})$. The shaded region is the unique irreducibility class $I$ of the Markov chain defined by the transition probabilities $R$, c.f. (50). Note that $\Delta\cup\Gamma\subset I$.
6 Conclusion
Discussion
In this work, we have explored the fundamental structure of random attractors for discrete-time random dynamical systems on discrete state spaces. In Section 3, we have shown the existence of a unique random attractor $A$, that is a forward attractor for compact random sets and a pullback attractor for finite deterministic sets. However, $A$ is typically not a pullback attractor for compact random sets. It is common to define pullback attraction for a class of suitable compact random sets, e.g. tempered random sets [KR11, Definition 14.3]. Investigating which classes of compact random sets are pullback attracted by $A$ — with or without assuming ergodicity of degree 2 — is a topic for potential future research.
Our investigations in Section 4 have structured the state space via insulated and synchronizing sets, showing that the $\mathds{P}$-a.s. size $\kappa$ of the attractor coincides with the size of any maximum insulated set. In all our examples of RDS representations of Markov chains, we have either found $\kappa=1$ or apparent symmetries yielding $\kappa>1$. In general, it is unclear whether $\kappa>1$ is only possible if the associated Markov chain admits some form of symmetry. The result that points in this direction is Proposition 4.7 (together with Remark 4.8), stating that $\kappa>1$ is only possible if the state space can be partitioned into elements of equal mass under the stationary distribution $\pi$.
Section 5 has characterized expected hitting times: we have shown that for any RDS representation of a Markov chain that is ergodic of degree 2, any initial condition hits the attractor in a time of finite mean. In particular, the expected time to hit the attractor, when the initial condition is distributed according to $\pi$, is finite as well (see formula (28)). An interesting open question is which RDS representation minimizes this quantity, i.e. for which RDS representation of a given Markov chain a randomly distributed initial conditions hits the attractor within the shortest time span. Identifying this RDS would be particularly useful for sampling from the stationary distribution $\pi$, using the method of the coupling from the past algorithm [PW96].
Outlook
The investigations developed in this paper presume a fixed RDS, including fixed parameters in the equations of motion. The aim of a bifurcation theory of RDS addresses the question on whether the dynamics of an RDS depending on a parameter $\lambda\in\mathbb{R}^{d}$ remains ‘equivalent’ whenever the parameter $\lambda$ undergoes slight variations. More specifically, a bifurcation occurs at some $\lambda_{0}$ if there are arbitrarily close values $\lambda$ leading to dynamics which are not topologically equivalent [Arn98]. A simple observation in the discrete state space setting is that the attractors of two topologically equivalent RDS have the same number of elements.
In certain situations more subtle distinctions in terms of topological equivalence become relevant. For instance, in [CDLR17], the notion of uniform topological equivalence is introduced to determine whether nearby systems preserve the (uniform) exponential attraction towards a random equilibrium. For discrete state spaces, the hitting time towards the attractor is finite almost surely, although its expected value may be finite or infinite. This qualitative difference suggests that a more refined notion of equivalence may distinguish between both cases.
On a related note, a change in the parameters of an RDS may also exhibit a qualitative change in the time series of the two-point motion. For instance, in chemical reaction networks, it was observed that the two-point motion of the Schlögl birth-death chain (as given by Gillespie’s algorithm) can exhibit a larger exploration of its state space upon variation of its kinetic constants [EOMUW24]. This phenomenon is intrinsically related to the bistability of the system. In terms of the dynamics of maximum insulated sets, this corresponds to an enlarging of the set $\mathcal{A}$, cf. (23). A thorough exploration of a bifurcation theory for attractors of RDS in discrete state spaces is left for future work.
Acknowledgements
This research has been partially supported by Deutsche Forschungsgemeinschaft (DFG) through grant CRC 1114 “Scaling Cascades in Complex Systems”, Project Number 235221301, Project A02 “Multiscale data and asymptotic model assimilation for atmospheric flows ” (G.O.-M.) and Project A08 “Characterization and prediction of quasi-stationary atmospheric states” (R.C. and M.E.), and through Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ EXC-2046/1, project 390685689, in particular subprojects AA1-8 (M.E. and G.O.-M.) and AA1-18 (M.E.). M.E. additionally thanks the Einstein Foundation for supporting his research. R.C. thanks Adam Schweitzer for engaging discussions, and providing valuable examples, like (21) in Example 1. The authors thank Francois Baccelli, Dennis Chemnitz and Tobias Hurth for helpful discussions and contributing valuable ideas.
Appendix A Assumption $\mathcal{F}=\mathcal{F}_{-\infty}^{\infty}$
Consider an irreducible and positive recurrent Markov chain on a countable state space $\mathds{X}$ with transition probabilities $P$ and unique stationary distribution $\pi$. Let $(\Omega,\mathcal{F},\mathds{P})$ be a probability space and $(\theta,\varphi)$ a Markov RDS representation of the Markov chain on $\mathds{X}$. Recall the definition of $\mathcal{F}_{-\infty}^{\infty}$ in (7). We state some additional results obtained under the following assumption.
Assumption 1.
$\mathcal{F}=\mathcal{F}_{-\infty}^{\infty}$.
This assumption asserts that $\mathcal{F}$ does not contain any information apart from the dynamical information of the cocycle $\varphi$. By definition, any RDS $(\theta,\varphi)$ is still a well-defined RDS over the probability space $(\Omega,\mathcal{F}_{-\infty}^{\infty},\mathds{P})$ with the same dynamical behavior. In particular, the dynamical properties of the RDS, like the attractor, insulated sets, synchronizing sets and the expected time until attraction, remain unaltered when equipping $\Omega$ with the $\sigma$-algebra $\mathcal{F}_{-\infty}^{\infty}$. However, Assumption 1 has some interesting consequences about the uniqueness of invariant measures and the indecomposability of the unique attractor. Before we state these results, we show that any RDS defined over a shift-space naturally satisfies Assumption 1.
Recall the definition of a shift-space $(\Sigma^{\mathds{Z}},\mathcal{G}^{\otimes Z},\mathds{P}_{0}^{\otimes\mathds{Z}},\theta)$ from Section 2.4, where $\Sigma$ is the set of maps $f:\mathds{X}\to\mathds{X}$. We consider the RDS over this shift-space defined by $\varphi_{\omega}(x)=\omega_{0}(x)$. We find
$\mathcal{F}_{-s}^{t}=\mathcal{G}^{\otimes\{-s,\ldots,t\}}.$ (54)
In particular, we have $\mathcal{F}_{-\infty}^{\infty}=\mathcal{G}^{\otimes Z}=\mathcal{F}$. Hence, any RDS $(\theta,\varphi)$ defined over a shift-space satisfies Assumption 1.
The correspondence theorem, cf. Theorem 2.7, asserts the existence of a unique invariant Markov measure $\mu$ defined by the sample measures $\mu_{\omega}=\lim_{n\to\infty}(\varphi^{n}_{\theta_{-n}\omega})_{*}\pi$. Under Assumption 1, the measure $\mu$ is ergodic with respect to the skew-product map $\Phi$, cf. [Arn98, Theorem 1.7.2]. In the following, we discuss the consequences of the ergodicity of $\mu$.
Proposition A.1.
$A$ is the only invariant compact random set.
Proof.
Assume there was another compact random invariant set $B$. [Arn98, Theorem 1.7.2] states that the unique invariant Markov measure $\mu$ is ergodic with respect to the skew-product map $\Phi$. Since $B$, as a subset of $\Omega\times\mathds{X}$, is invariant under $\Phi$, we have $\mu(B)\in\{0,1\}$. Let $\overline{\Omega}\subset\Omega$ be a $\theta$-invariant set of full measure on which $B(\omega)$ is finite and $\varphi$-invariant. Note that
$|B(\theta\omega)|=|\varphi_{\omega}(B(\omega))|\leq|B(\omega)|.$
Hence, the function $\omega\mapsto|B(\omega)|$ is non-increasing under $\theta$ and measurable. By ergodicity of $\theta$, there is $\kappa^{\prime}\in\mathds{N}$ such that $|B(\omega)|=\kappa^{\prime}$ for all $\omega\in\overline{\Omega}$.
Since $A$ is a weak attractor for compact random sets, we find
$\lim_{n\to\infty}\mathds{P}(\varphi_{\omega}^{n}(B(\omega))\subset A(\theta_{n}\omega))$ $=\lim_{n\to\infty}\mathds{P}(B(\theta_{n}\omega)\subset A(\theta_{n}\omega))$ (55)
$=\mathds{P}(B(\omega)\subset A(\omega))$
$=1.$
Hence, $B(\omega)$ is $\mathds{P}$-a.s. contained in $A(\omega)$. Assuming that $B$ is different from $A$, we get $1\leq\kappa^{\prime}<\kappa$, where $\kappa$ is the $\mathds{P}$-a.s. size of $A(\omega)$. We compute the $\mu$-measure of $B$, cf. (14),
$\mu(B)$ $=\mathds{E}\left[\sum_{x\in\mathds{X}}\mu_{\omega}(x)\mathds{1}_{B(\omega)}(x)\right]$ (56)
$=\mathds{E}\left[\sum_{a\in A(\omega)}\frac{1}{\kappa}\mathds{1}_{B(\omega)}(a)\right]$
$=\frac{\kappa^{\prime}}{\kappa}\in(0,1).$
This contradicts $\mu(B)\in\{0,1\}$ such that the claim follows. ∎
Corollary A.2.
There are no other invariant measures on $\Omega\times\mathds{X}$ apart from the unique invariant Markov measure $\mu$.
Proof.
Consider an invariant measure $\nu$. Since $A$ is a forward attractor for compact random sets, the support of $\nu_{\omega}$ has to be contained in $A(\omega)$ for almost all $\omega\in\Omega$, implying $|\text{supp}(\nu_{\omega})|\leq\kappa=|A(\omega)|$. Assume $B(\omega):=\{x\in\mathds{X}\>|\>\nu_{\omega}(x)>\frac{1}{\kappa}\}\neq\emptyset$. Since $\nu$ is invariant, this defines an invariant compact random set. By Proposition A.1, we have that $B(\omega)=A(\omega)$ for almost all $\omega\in\Omega$. This is a contradiction, since $|B(\omega)|<\kappa$ by construction. Hence, $\text{supp}(\nu_{\omega})=A(\omega)$ and $\nu_{\omega}(x)=1/\kappa$ for $x\in A(\omega)$, i.e. $\nu=\mu$ by the characterization (14) of the Markov measure $\mu$. ∎
We conclude this subsection with an example of an RDS $(\theta,\varphi)$ that does not satisfy Assumption 1 and admits non-Markovian invariant measures.
Example 5.
Consider the Markov chain on a two-element space $\mathds{X}=\{A,B\}$ with transition probabilities $P(A,B)=P(B,A)=1$. This Markov chain is irreducible and positive recurrent with unique stationary distribution $\pi=(\frac{1}{2},\frac{1}{2})$. Define a probability space by $\Omega=\{\omega_{1},\omega_{2}\}$, equipped with the power set $\sigma$-algebra $\mathcal{F}=\mathcal{P}(\Omega)$, and define the probability measure $\mathds{P}(\omega_{1})=\mathds{P}(\omega_{2})=\frac{1}{2}$. Define a measure-preserving transformation $\theta:\Omega\to\Omega$ by $\theta(\omega_{1})=\omega_{2}$, and $\theta(\omega_{2})=\omega_{1}$. The cocycle $\varphi$ is deterministic and given by $\varphi_{\omega}(A)=B$ and $\varphi_{\omega}(B)=A$, for all $\omega\in\Omega$. One can check that $(\theta,\varphi)$ is a well-defined RDS which represents the Markov chain on $\mathds{X}$. Note that =$\mathcal{F}_{-\infty}^{\infty}=\{\emptyset,\Omega\}\neq\mathcal{P}(\Omega)$ such that Assumption 1 is not satisfied. The unique invariant Markov measure on $\Omega\times\mathds{X}$, which is given by the correspondence theorem, satisfies $\mu(\omega,x)=\frac{1}{4}$, for each $(\omega,x)\in\Omega\times\mathds{X}$. However, the skew-product map $\Phi$ has two invariant components, namely $\{(A,\omega_{1}),(B,\omega_{2})\}$ and $\{(A,\omega_{2}),(B,\omega_{1})\}$, each of which is $\mathcal{F}\otimes\mathcal{P}(\mathds{X})$-measurable and has $\mu$-measure $\frac{1}{2}$. Hence, $\mu$ is not ergodic with respect to $\Phi$. The restriction of $\mu$ to the two $\Phi$-invariant subsets yields two invariant non-Markovian measures which are ergodic with respect to $\Phi$. $\Diamond$
Appendix B Petite sets and Lyapunov functions
We formally define the notion of a petite set of a Markov chain and introduce a Foster–Lyapunov type drift condition. In this brief introduction, we closely follow [MT12], which considers general state spaces.
First, we introduce the notion of $\psi$-irreducibility, which is a generalized concept of irreducibility for uncountable state spaces.
Definition B.1.
A Markov chain $(X_{n})_{n\in\mathds{N}}$ on a state space $\mathds{X}$ is called $\psi$-irreducible if there is a measure $\psi$ on $\mathds{X}$ such that $\mathds{P}_{x}[\tau_{A}<\infty]=1$ for all $x\in\mathds{X}$ and sets $A\subset\mathds{X}$ with $\psi(A)>0$.
In the case that $\mathds{X}$ is countable, a measure $\psi$ is simply a function $\psi:\mathds{X}\to\mathds{R}_{\geq 0}$. Hence, if $I\subset\mathds{X}$ is a irreducibility class of $(X_{n})_{n\in\mathds{N}}$, and $\mathds{P}_{x}(\tau_{I}<\infty)=1$ for all $x\in\mathds{X}$, then the Markov chain is $\psi$-irreducible with $\psi=\mathds{1}_{I}$. Conversely, if $(X_{n})_{n\in\mathds{N}}$ is $\psi$-irreducible, then $I:=\text{supp}(\psi)$ is the unique irreducibility of the Markov chain and $\mathds{P}_{x}(\tau_{I}<\infty)=1$ for all $x\in\mathds{X}$. Hence, a Markov chain on a countable state space is $\psi$-irreducible if and only if there is a unique irreducibility class that is reached $\mathds{P}_{x}$-a.s. for any $x\in\mathds{X}$.
Next, we introduce petite sets. Let $(X_{n})_{n\in\mathds{N}}$ be a Markov chain on a countable state space $\mathds{X}$ with transition probabilities $P$. Let $(a_{n})_{n\in\mathds{N}}$ be a probability distribution on $\mathds{N}$, i.e. $a_{n}\geq 0$ and $\sum_{n\in\mathds{N}}a_{n}=1$. We define the sampled transition probabilities
$Q_{a}(x,y):=\sum_{n\in\mathds{N}}a_{n}P^{n}(x,y).$ (57)
These transition probabilities correspond to sampling the Markov chain $(X_{n})_{n\in\mathds{N}}$ at i.i.d. time intervals which are distributed according to $(a_{n})_{n\in\mathds{N}}$.
Definition B.2.
A set $C\subset\mathds{X}$ is called petite, if there is a sampling distribution $(a_{n})_{n\in\mathds{N}}$ and a non-trivial measure $\nu_{a}$ on $\mathds{X}$ such that
$Q_{a}(x,y)\geq\nu_{a}(y),$ (58)
for all $x\in C$ and $y\in\mathds{X}$.
A proof of the following result is found in [MT12, Proposition 5.5.5].
Proposition B.3.
Assume that $(X_{n})_{n\in\mathds{N}}$ is $\psi$-irreducible. Then, the union of two petite sets is petite.
Finally, we introduce Lyapunov functions and a Foster–Lyapunov type drift condition. Let $(X_{n})_{n\in\mathds{N}}$ be a Markov chain on a countable state space $\mathds{X}$ with transition probabilities $P$. A Lyapunov function is a map $V:\mathds{X}\to\mathds{R}_{\geq 0}$ that assigns an ’energy’ to each state $x\in\mathds{X}$. The expected energy after one step of the Markov chain is given by
$PV(y)=\sum_{x\in\mathds{X}}P(x,y)V(y).$ (59)
Let $C\subset\mathds{X}$. We say that $V$ satisfies a drift condition towards $C$ if there is a constant $b\geq 0$ such that
$PV(x)\leq V(x)-1+b\mathds{1}_{C}(x),$ (60)
for all $x\in\mathds{X}$. The following result is found in [MT12, Theorem 11.3.11].
Theorem B.4.
Suppose that $(X_{n})_{n\in\mathds{N}}$ is a $\psi$-irreducible Markov chain on a countable state space $\mathds{X}$. Assume there is a function $V:\mathds{X}\to\mathds{R}_{\geq 0}$ that satisfies (60) with a constant $b\geq 0$ and a petite set $C$. Then, for each $y\in\text{supp}(\psi)$, there is a constant $c(y)$ such that
$\mathds{E}_{x}[\tau_{y}]\leq V(x)+c(y).$ (61)